### v4

# Exact Solution to Data-Driven Inverse Optimization of MILPs in Finite Time via Gradient-Based Methods

Abstract.
A data-driven inverse optimization problem (DDIOP) seeks to estimate an objective function (i.e., weights) that is consistent with observed optimal-solution data, and is important in many applications, including those involving mixed integer linear programs (MILPs). In the DDIOP for MILPs, the prediction loss on features (PLF), defined as the discrepancy between observed and predicted feature values, becomes discontinuous with respect to the weights, which makes it difficult to apply gradient-based optimization. To address this issue, we focus on a Lipschitz continuous and convex suboptimality loss. By exploiting its convex and piecewise-linear structure and the interiority of the minimum set, we show that a broad class of gradient-based optimization methods, including projected subgradient descent (PSGD), reaches the minimum suboptimality loss value in a finite number of iterations, thereby exactly solving the DDIOP for MILPs. Furthermore, as a corollary, we show that PSGD attains the minimum PLF in finitely many iterations. We also derive an upper bound on the number of iterations required for PSGD to reach finite convergence, and confirm the finite-step behavior through numerical experiments.
Key words and phrases:
inverse optimization problem, mathematical programming, mixed integer linear programming, projected subgradient method
AMS subject classifications 2020: 90C90(primary), 90C25, 90C52, 90C11, 90C05, 68Q25 (secondary)
1. Introduction
Many diverse processes and systems, ranging from human decision-making to natural phenomena, can often be described by mathematical optimization models. However, in practice, the true objective function underlying such models is frequently unknown a priori. Consequently, the inverse optimization problem—which aims to estimate the objective function from observed data (Ahuja and Orlin, 2001; Heuberger, 2004; Chan et al., 2019, 2023)—is of substantial practical importance.
Inverse optimization has been extensively studied (Ahuja and Orlin, 2001; Heuberger, 2004; Chan et al., 2019, 2023) and applied in a broad range of domains, including geophysics (Tarantola, 2005; Burton and Toint, 1992), transportation (Bertsimas et al., 2015), power systems (Birge et al., 2017), television advertisement scheduling (Suzuki et al., 2019), and healthcare (Chan et al., 2022). Moreover, it has also developed as a foundational framework for machine learning methods, including inverse reinforcement learning (Ng et al., 2000) and contrastive learning (Shi et al., 2023).
In this paper, we consider a fundamental and important setting in which the objective function of the forward optimization problem is represented as a linear combination of multiple functions (cf. Murata et al., 1996; Konak et al., 2006; Gunantara, 2018), with particular emphasis on settings that include mixed integer linear programs (MILPs). Let $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$ be the decision space. For each state $s\in\mathcal{S}$, let $X(s)\subset\mathcal{X}$ be a bounded and closed feasible region. For each $i=1,\ldots,d$, let $f_{i}\colon\mathcal{X}\to\mathbb{R}$ be a piecewise linear mapping, and define $f:=(f_{1},\ldots,f_{d})$. Let $\Theta\subset\mathbb{R}^{d}$ be a closed convex set representing the weight space, and let $\theta=(\theta_{1},\ldots,\theta_{d})\in\Theta$ denote the weights. We define the optimal solution to the forward problem and its associated feature vector as follows:
$x^{*}(\theta,s)\in\operatorname*{arg\,max}_{x\in X(s)}\theta^{\top}f(x)=\sum_{i=1}^{d}\theta_{i}f_{i}(x),\quad a^{*}(\theta,s)=f(x^{*}(\theta,s)).$ (1.1)
Let $\theta^{*}\in\Theta$ denote the (unknown) true weights that generated the observed data. For each $n=1,\ldots,N$, suppose that we are given $s^{(n)}\in\mathcal{S}$ and the observed optimal solution $x^{(n)}=x^{*}(\theta^{*},s^{(n)})$. The data-driven inverse optimization problem (DDIOP) associated with the forward optimization problem Equation˜1.1 is defined as the problem of finding $\theta\in\Theta$ satisfying the following condition for all $n=1,\ldots,N$:
$x^{(n)}\in\operatorname*{arg\,max}_{x\in X(s^{(n)})}\theta^{\top}f(x)=\sum_{i=1}^{d}\theta_{i}f_{i}(x).$ (1.2)
An illustration of solving the DDIOP for MILPs is shown in Figure˜1.
Objective Function $f_{i}$ Coefficient (cost/time) $\theta_{i}$
Tardiness of job 11 ???
$\vdots$ $\vdots$
Tardiness of job $D$ ???
Input 1
Objective Function $f_{i}$ Coefficient (cost/time) $\theta_{i}$
Tardiness of job 11 0.15
$\vdots$ $\vdots$
Tardiness of job $D$ 0.31
OutputInverse OptimizationInput 2: Schedules
Figure 1. Illustration of the DDIOP. This figure depicts the estimation of the objective-function coefficients such that the given shift schedule becomes an optimal solution.
As evaluation criteria for whether Equation˜1.2 has been solved, one may consider the prediction loss of features (PLF) (Aswani et al., 2018; Chan et al., 2019; Babier et al., 2021; Chan et al., 2023; Ferber et al., 2023; Liang and Dvorkin, 2023) and the suboptimality loss (Mohajerin Esfahani et al., 2018). The PLF, denoted by $\ell_{\mathrm{plf}}(\theta)$, is given by $\ell_{\mathrm{plf}}(\theta)=\sum_{n=1}^{N}\left\lVert a^{*}(\theta,s^{(n)})-a^{(n)}\right\rVert_{2}^{2}/N,$ where $a^{(n)}=f(x^{(n)})$. If the PLF equals 0, then Equation˜1.2 is solved.
However, in the DDIOP for MILPs, the PLF is discontinuous with respect to $\theta\in\Theta$. As a consequence, gradient-based optimization methods are not directly applicable (cf. Beck, 2017; Hazan, 2019; Garrigos and Gower, 2023). As a naive approach to solving Equation˜1.2, one may discretize the weight space $\Theta=\Delta^{d-1}$ by taking a uniform set of points (uniform point approximation; UPA), or by sampling points at random (random point approximation; RPA). However, letting $T$ denote the number of points, the distance between the true weight $\theta^{*}$ and the learned point set is on the order of $O(T^{-1/(d-1)})$ for UPA (Flatto and Newman, 1977, Theorem 2.1) and $O((\log T/T)^{1/(d-1)})$ for RPA (Reznikov and Saff, 2015, Corollary 2.3). These rates deteriorate rapidly as the dimension $d$ increases.
On the other hand, as a Lipschitz continuous and convex loss function for assessing whether the DDIOP for MILPs (Equation˜1.2) has been solved, one may consider the suboptimality loss
$\ell_{\mathrm{sub}}(\theta):=\frac{1}{N}\sum_{n=1}^{N}\left(\theta^{\top}a^{*}(\theta,s^{(n)})-\theta^{\top}a^{(n)}\right),$
as proposed in (Mohajerin Esfahani et al., 2018). Solving Equation˜1.2 is equivalent to finding $\theta$ such that $\ell_{\mathrm{sub}}(\theta)=0$. Therefore, by minimizing the suboptimality loss, one may expect to obtain a solution to Equation˜1.2.
For minimizing the suboptimality loss, a variety of standard first-order (online) optimization methods have been used in the inverse optimization literature (Bärmann et al., 2018; Besbes et al., 2021, 2025; Gollapudi et al., 2021; Kitaoka and Eto, 2023a, b; Sakaue et al., 2025). Typical examples include projected subgradient descent (PSGD) (Boyd et al., 2003; Beck, 2017), multiplicative weights update (MWU) (Arora et al., 2012), online Newton step (ONS) (Hazan et al., 2007), and MetaGrad (van Erven and Koolen, 2016; van Erven et al., 2021). These methods come with well-established regret/convergence guarantees for Lipschitz convex losses, which imply that $\ell_{\mathrm{sub}}$ decreases as the number of iterations $T$ grows.
However, much of the existing literature primarily provides asymptotic guarantees, i.e., $\ell_{\mathrm{sub}}(\theta^{t})$ $\to 0$ only in the limit as $T\to\infty$. In contrast, whether one can exactly solve the DDIOP for MILPs within finitely many iterations—namely, guarantee $\ell_{\mathrm{sub}}=0$ after a finite number of updates—has not been sufficiently clarified.
Contributions
This paper establishes that, for the DDIOP Equation˜1.2 including MILPs, gradient-based optimization methods for minimizing the suboptimality loss $\ell_{\mathrm{sub}}$ reach $\ell_{\mathrm{sub}}=0$ in finitely many iterations, and provides a framework for evaluating the required number of iterations. Here, by a “gradient-based optimization method” we mean an iterative procedure that updates $\theta$ by using (sub)gradient information of $\ell_{\mathrm{sub}}$. Specifically, we show the following.
  • (Finite-step exact solvability). If a gradient-based optimization method (A) can be described solely in terms of the past iterates and subgradients, and (B) enjoys a standard convergence guarantee for arbitrary Lipschitz convex functions, then it can reduce the suboptimality loss for MILPs to 0 in finitely many iterations; equivalently, it can exactly solve the DDIOP Equation˜1.2 (Theorem˜4.3).
  • (Explicit iteration complexity). By applying Theorem˜4.3 to PSGD with representative step-size rules, we derive an upper bound on the number of iterations required for finite-step attainment, expressed in terms of a constant $\gamma(\ell_{\mathrm{sub}})$ (Equation˜4.3) (˜4.2 and 4.2). A comparison with existing results is summarized in Table˜1.
  • (Key proof ingredients: convexity, piecewise linearity, and interior-point property). The key structural properties are that the suboptimality loss is convex and piecewise linear, and that (for almost every true weight $\theta^{*}\in\Theta$) the set of minimizers has a nonempty interior. Under these conditions, standard convergence guarantees for gradient-based optimization methods can be strengthened—via the gap parameter $\gamma(\ell_{\mathrm{sub}})$—to yield finite-step convergence to the minimum. An intuitive explanation is provided in ˜5.
Table 1. Comparison of approaches for solving the DDIOP Equation˜1.2 for MILPs (suboptimality loss). Whereas prior work primarily provides asymptotic guarantees—i.e., the loss decreases as $T\to\infty$—the present paper differs in that it guarantees $\ell_{\mathrm{sub}}=0$ within finitely many iterations. For a detailed comparison (explicit rates and constants), see Table˜4.
Method (representative examples) Typical guarantee (informal) $\ell_{\mathrm{sub}}=0$
Point-set search (UPA/RPA) Approximation based on covering error (strong dimension dependence) No
Online optimization (MWU / ONS / MetaGrad, etc.) Asymptotic decrease of $\ell_{\mathrm{sub}}$ (e.g., $O(1/\sqrt{T})$, $O(\log T/T)$) No
This paper (gradient-based) Finite-step attainment of 0 by leveraging the structure of $\ell_{\mathrm{sub}}$ (convex, piecewise linear, plus an interior-point property) Yes
As shown in Table˜1, while existing approaches provide asymptotic error bounds for $\ell_{\mathrm{sub}}$, they do not, in general, guarantee that $\ell_{\mathrm{sub}}=0$ is attained at some finite iteration $T$. This paper shows that, for gradient-based optimization methods including PSGD, the DDIOP for MILPs can be solved exactly in finitely many iterations. Moreover, for specific standard choices of PSGD parameters, we provide an upper bound on the number of iterations required to reach $\ell_{\mathrm{sub}}=0$ (Theorem˜4.3 and ˜4.2 and 4.2).
Analogously to the discussion on DDIOP, we show that, also for PLFs associated with MILPs, PSGD applied to the suboptimality loss attains the minimum value of the PLF in a finite number of iterations (Theorems˜4.10 and 4.11). A comparison between existing approaches and the proposed method for PLFs is presented in Table˜5.
In the numerical experiments, using synthetic data and a scheduling problem, we empirically demonstrate that gradient-based optimization methods can reach the minimum value of the PLF within finitely many iterations. This provides empirical evidence that the DDIOP for MILPs can be solved in a finite number of iterations.
The remainder of this paper is organized as follows. In ˜2, we review related work. In ˜3, we describe the DDIOP setting and present our proposed algorithm, which applies gradient-based optimization to the suboptimality loss. In ˜4, we show that the proposed method solves the DDIOP for MILPs within finitely many iterations and provide an upper bound on the required number of iterations. We further show that, when PSGD is instantiated within the proposed framework, the PLF can be driven to 0 in finitely many iterations, together with an upper bound on the iteration complexity. In ˜5, we provide an intuitive explanation of the proofs of the main theorems presented in ˜4. In ˜6, we report numerical experiments, demonstrating that the proposed method reaches the minimum value of the PLF within finitely many iterations. Finally, ˜7 concludes the paper and discusses directions for future work.
2. Related Work
Methods for smoothing the prediction loss of optimal solutions.
Let the weight space be $\Theta=\Delta^{d-1}$. To address the difficulty that the PLF takes only discrete values as a function of $\theta\in\Delta^{d-1}$, there exist approaches other than our method, including adding noise to the inputs (Berthet et al., 2020) and adding a regularization term to the objective (Wilder et al., 2019). However, both Wilder et al. (2019) and Berthet et al. (2020) suffer from several issues regarding their theoretical guarantees; see ˜B.1.
Online optimization of the suboptimality loss.
As an approach to offline optimization of the suboptimality loss, one may apply online optimization methods. By leveraging regret analysis—the standard performance measure in online optimization—one can obtain guarantees for the corresponding offline optimization problem. For existing regret analyses for the suboptimality loss and their results (Bärmann et al., 2018; Besbes et al., 2021, 2025; Gollapudi et al., 2021; Sakaue et al., 2025), see Appendix˜E. For the offline implications derived from regret guarantees, see Table˜1.
Relation to NP-hardness results.
Our finite-time guarantee is expressed in terms of the gap parameter $\gamma(\ell_{\mathrm{sub}})$. For instance, for PSGD with square-root step sizes, the resulting upper bound typically scales as $T=O(1/\gamma(\ell_{\mathrm{sub}})^{2})$ up to polynomial factors in $L(\ell_{\mathrm{sub}})$ and $\mathrm{diam}(\Theta)$. Therefore, our results do not imply a polynomial-time algorithm for PLF minimization in general. In fact, the NP-hardness result of Aswani et al. (2018, Theorem 1) suggests that, as the problem size grows (e.g., the number of constraints and the dimension increase), $\gamma(\ell_{\mathrm{sub}})$ need not admit a polynomial lower bound and can become exponentially small, which would make the above iteration bound exponentially large.
For additional related works, see Appendix˜B.
3. Problem Setup and Algorithms
We impose the following assumptions to specify the problem setting studied in this paper.
Assumption 3.1.
Let the nonempty set $\mathcal{S}$ be the state space, and let the nonempty set $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$ be the set of decision variables. Let $f=(f_{1},\ldots,f_{d})\colon\mathcal{X}\to\mathbb{R}^{d}$ be a mapping such that each component $f_{i}$ is a piecewise linear function. Let the weight space $\Theta\subset\mathbb{R}^{d}$ be a bounded, closed, and convex set. For each state $s\in\mathcal{S}$, let the feasible region $X(s)$ be a finite union of bounded, closed, and convex polytopes that are subsets of $\mathcal{X}$. Let $\mathcal{D}:=\left\{\left(s^{(n)},x^{(n)}\right)\right\}_{n=1}^{N}$ be a set of training data such that the samples $s^{(n)}\in\mathcal{S}$ are generated from an unknown distribution $\mathbb{P}_{\mathcal{S}}$, and there exists an unknown $\theta^{*}\in\Theta$ satisfying, for each $n=1,\ldots,N$, $x^{(n)}=x^{*}(\theta^{*},s^{(n)})$.
As a technical condition, we further assume the following.
Assumption 3.2.
For each $n=1,\ldots,N$, the feature $a^{*}(\theta^{*},s^{(n)})$ is uniquely determined.
˜3 is a natural assumption, in view of the following statement.
Lemma 3.3.
Suppose ˜3 holds and let $\Theta=\Delta^{d-1}\subset\mathbb{R}^{d}$. Then, for $\theta\in\Delta^{d-1}$ almost everywhere (with respect to the measure on $\Delta^{d-1}$ induced by the Lebesgue measure), the feature $a^{*}(\theta,s^{(n)})$ is uniquely determined for every $n=1,\ldots,N$.
The proof of this lemma is given in Appendix˜I.
In this paper, based on the following proposition, we apply a gradient-based optimization method to the suboptimality loss $\ell_{\mathrm{sub}}$ and minimize it.
Proposition 3.4 (Bärmann et al. 2018, Proposition 3.1; Kitaoka and Eto 2023a, Lemma 4.8).
Suppose that ˜3 holds. Then the following statements hold: (A) the suboptimality loss $\ell_{\mathrm{sub}}$ is convex; (B) the suboptimality loss $\ell_{\mathrm{sub}}$ is Lipschitz continuous; and (C) a subgradient of the suboptimality loss $\ell_{\mathrm{sub}}$ at $\theta\in\Theta$ is given by $g(\theta):=\sum_{n=1}^{N}\bigl(a^{*}(\theta,s^{(n)})-a^{(n)}\bigr)/N$.
Algorithm 1 Minimization of the suboptimality loss
1: Initialize $\theta^{1}\in\Delta^{d-1}$
2:for $t=1,\ldots,T-1$ do
3:  For each $n=1,\ldots,N$, solve for $x^{*}(\theta^{t},s^{(n)})$
4:  $\theta^{t+1}\leftarrow\mathrm{update}_{t}\!\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t}\mid\ell_{\mathrm{sub}},g\right)$
5:end for
6:return $\theta_{\mathrm{best}}^{T}\in\operatorname*{arg\,min}_{\theta\in\{\theta^{t}\}_{t=1}^{T}}\ell_{\mathrm{sub}}(\theta)$
Hereafter, when it is clear from the context, $\min_{\Theta}\ell=\min_{\theta\in\Theta}\ell(\theta)$ and $\operatorname*{arg\,min}_{\Theta}\ell=\operatorname*{arg\,min}_{\theta\in\Theta}\ell(\theta)$ are used as abbreviations. Let $\ell\colon\Theta\to\mathbb{R}$ be a Lipschitz-continuous convex function, and let $\partial\ell$ denote the subdifferential of $\ell$. Let $\nabla\ell\colon\Theta\to\mathbb{R}^{d}$ be any (sub)gradient selection satisfying, for every $\theta$, $\nabla\ell(\theta)\in\partial\ell(\theta)$. For $t\in\mathbb{Z}_{\geq 1}$, let the update map be $\mathrm{update}_{t}\colon\Theta^{t}\times\mathbb{R}^{t}\times(\mathbb{R}^{d})^{t}\to\Theta$. Moreover, define
$\begin{split}&\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t}\mid\ell,\nabla\ell\right)\\ &:=\mathrm{update}_{t}(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t},\{\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t},\{\nabla\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t})\end{split}$
We refer to the sequence $\{\mathrm{update}_{t}\}_{t}$ as an update rule. The resulting procedure is described in Algorithm˜1.
Example 3.5 (Projected subgradient descent).
Let $\alpha_{t}\colon\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a mapping, and refer to $\{\alpha_{t}\}_{t\in\mathbb{Z}_{\geq 1}}$ as the learning rate. Let $\operatorname{Proj}_{\Theta}$ denote the projection onto $\Theta$. We identify the mapping $\alpha_{t}$ with its value $\alpha_{t}(\theta^{t},\ell(\theta^{t}),\nabla\ell(\theta^{t}))$ when no confusion arises. We define projected subgradient descent (PSGD) (cf. Boyd et al., 2003; Beck, 2017) as the update rule $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 by $\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t}|\ell,\nabla\ell\right)=\operatorname{Proj}_{\Theta}\left(\theta^{t}-\alpha_{t}\nabla\ell(\theta^{t})\right).$ The learning rate $\alpha_{t}$ is said to be a nonsummable step size (NSS) if there exists a sequence $\{\beta_{t}\}_{t}\subset\mathbb{R}$ such that
$\beta_{t}>0,\quad\lim_{t\to\infty}\frac{\sum_{t=1}^{\infty}\beta_{t}^{2}}{\sum_{t=1}^{\infty}\beta_{t}}=0$ (3.1)
and $\alpha_{t}=\beta_{t}$. The learning rate $\alpha_{t}$ is said to be a square root step size (SRSS) if, for some $\beta>0$, it is an NSS learning rate with $\beta_{t}=\beta t^{-1/2}$. The learning rate $\alpha_{t}$ is said to be a nonsummable step length (NSL) if, for some sequence $\{\beta_{t}\}_{t}\subset\mathbb{R}$ satisfying Equation˜3.1, we set $\alpha_{t}=\beta_{t}\|\nabla\ell(\theta^{t})\|^{-1}$ when $\nabla\ell(\theta^{t})\not=0$, and $\alpha_{t}=0$ when $\nabla\ell(\theta^{t})=0$. The learning rate $\alpha_{t}$ is said to be a square root step length (SRSL) if, for some $\beta>0$, it is an NSL learning rate with $\beta_{t}=\beta t^{-1/2}$. The learning rate $\alpha_{t}$ is said to be Polyak if the minimum value $\min_{\Theta}\ell$ is known and $\alpha_{t}=\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)\|\nabla\ell(\theta^{t})\|^{-2}$ holds.
We present below an example satisfying ˜3, together with an implementation example of Algorithm˜1.
Example 3.6.
In ˜3, let $f=\operatorname{Id}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$. Under this setting, Equation˜1.1 becomes an MILP. Let the sequence $\{\mathrm{update}_{t}\}_{t}$ be PSGD, and implement the projection $\operatorname{Proj}_{\Delta^{d-1}}$ onto $\Theta=\Delta^{d-1}$ as in Wang and Carreira-Perpinán (2013). Then one can implement Algorithm˜1. This implementation is identical to Kitaoka and Eto (2023a, Algorithm 1).
Remark 3.7.
In ˜3, it is more natural to require that the weight space $\Theta$ does not contain 0; see Appendix˜H for the rationale.
4. Main Results
4.1. Finite-time exact solvability of DDIOP for MILP (gradient based optimization method)
For gradient-based optimization methods, we make the following assumptions: namely, that the update rule depends only on the past sequence of iterates and (sub)gradients, and that a convergence rate in terms of the best iterate is guaranteed, respectively.
Assumption 4.1.
For any $t\in\mathbb{Z}_{\geq 1}$, we assume that $\mathrm{update}_{t}\colon\Theta^{t}\times\mathbb{R}^{t}\times(\mathbb{R}^{d})^{t}\to\Theta$ does not depend on the second component space $\mathbb{R}^{t}$.
Assumption 4.2.
For any $L$-Lipschitz-continuous convex function $\ell\colon\Theta\to\mathbb{R}$, any initial point $\theta^{1}\in\Theta$, and any $t\in\mathbb{Z}_{\geq 1}$, define the iterates inductively by $\theta^{t+1}=\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t}\mid\ell,\nabla\ell\right)$. Then there exists a nonincreasing function $Q_{L,\theta^{1}}\colon\mathbb{R}_{>0}\to\mathbb{R}$ such that, for every $\varepsilon>0$, if $T\geq Q_{L,\theta^{1}}(\varepsilon)$, then for any $L$-Lipschitz-continuous convex function $\ell\colon\Theta\to\mathbb{R}$, the best iterate is $\varepsilon$-accurate, that is, $\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\leq\varepsilon$.
Among PSGD methods, those with NSS or NSL learning rates satisfy ˜4.1, and explicit convergence rates for the best iterate are available (˜J.1 and J.2). In contrast, for PSGD with the Polyak learning rate, $\mathrm{update}_{t}$ depends on the second argument, i.e., the sequence of objective values $\{\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t}$, and hence it does not satisfy ˜4.1.
Since $X(s^{(n)})$ is a finite union of polyhedra and $f$ is such that each component function $f_{i}$ is piecewise affine, it follows that $f\!\left(X(s^{(n)})\right)$ is also a finite union of polyhedra (see Appendixes˜K and K). Let $Y^{(n)}$ denote the set of vertices of the polyhedral complex $f\!\left(X(s^{(n)})\right)$ (for the definition of the vertex set, see Appendix˜K). Under ˜3, by the maximum principle,
$a^{*}(\theta^{*},s^{(n)})\in Y^{(n)}$ (4.1)
holds. We define the Lipschitz constant $L(\ell_{\mathrm{sub}})$ and a positive constant $\gamma(\ell_{\mathrm{sub}})$ as follows:
$L(\ell_{\mathrm{sub}})$ $:=\sup_{\begin{subarray}{c}a^{n}\in f(X(s^{(n)})),\\ n=1,\ldots,N\end{subarray}}\left\|\frac{1}{N}\sum_{n=1}^{N}(a^{n}-a^{(n)})\right\|\left(\leq L(f)\mathrm{diam}(\mathcal{X})\right),$ (4.2)
$\gamma(\ell_{\mathrm{sub}})$ $:=\max_{\theta\in\Theta}\min_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{\top}\left(a^{(n)}-a^{n}\right).$ (4.3)
Here, $L(f)$ denotes the Lipschitz constant of $f$, and $\mathrm{diam}(\mathcal{X})=\sup_{x,x^{\prime}\in\mathcal{X}}\|x-x^{\prime}\|$ denotes the diameter of $\mathcal{X}$. For an intuitive explanation of the constant $\gamma(\ell_{\mathrm{sub}})$, see ˜5. For the fact that $\gamma(\ell_{\mathrm{sub}})$ is positive, see ˜L.1 and L.2.
Our main theorem is as follows.
Theorem 4.3.
Suppose that ˜3, 4.1, 4.1 and 3 hold. Then,
if $T\geq Q_{L(\ell_{\mathrm{sub}}),\theta^{1}}(\gamma(\ell_{\mathrm{sub}}))$, we have $\min_{t=1,\ldots,T}\ell_{\mathrm{sub}}(\theta^{t})=\min_{\Theta}\ell_{\mathrm{sub}}=0$.
See Appendix˜L for the proof. From ˜3 and Theorem˜4.3, we obtain the following corollary.
Corollary 4.4.
Suppose that ˜3, 4.1 and 4.1 hold. Let $\Theta=\Delta^{d-1}$. Then, for Lebesgue-almost every $\theta^{*}\in\Delta^{d-1}$ (with respect to the measure on $\Delta^{d-1}$ induced by Lebesgue measure), the following holds: if $T\geq Q_{L(\ell_{\mathrm{sub}}),\theta^{1}}(\gamma(\ell_{\mathrm{sub}}))$, then $\min_{t=1,\ldots,T}\ell_{\mathrm{sub}}(\theta^{t})=\min_{\Theta}\ell_{\mathrm{sub}}=0$.
proof.
By ˜3, ˜3 holds for a.e. $\theta^{*}\in\Delta^{d-1}$. Fix such a $\theta^{*}$. Then Theorem˜4.3 is applicable, which yields the claim. ∎
Remark 4.5.
By strengthening the result of Theorem˜4.3, one can, in the case $\Theta=\Delta^{d-1}$, replace the constant $\gamma(\ell_{\mathrm{sub}})$ by a constant depending only on $\{X(s^{(n)})\}_{n}$, $f$, and $\Theta$ (see Theorem˜M.3).
4.2. Concrete complexity bounds for PSGD
In this section, by applying Theorem˜4.3, we bound the number of iterations required to solve the DDIOP associated with an MILP when incorporating a concrete gradient-based optimization method into the update rule of Algorithm˜1.
Let $\mathrm{diam}(\Theta)$ denote the diameter of $\Theta$. In what follows, we use the fact that each variant of PSGD satisfies ˜4.1 and admits an error bound for the best iterate (for NSS, see ˜J.1; for NSL, see ˜J.2). Then, by Theorem˜4.3, the following results hold.
PSGD (NSS)
Corollary 4.6.
Under ˜3 and 3, Algorithm˜1 with PSGD (NSS) solves Equation˜1.2 exactly in finitely many iterations.
When the learning rate is SRSS, one can explicitly construct the function $Q_{L,\theta^{1}}$ (see ˜J.1); hence, one can bound the number of iterations required to solve the DDIOP associated with an MILP as follows:
Corollary 4.7.
Under ˜3 and 3, Algorithm˜1 with PSGD (SRSS) solves Equation˜1.2 exactly within at most
$\max\left(1,\left(\frac{\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}L(\ell_{\mathrm{sub}})^{2}}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right)$ iterations.
For the construction of $Q_{L,\theta^{1}}$ for NSS (including SRSS) and the proof of ˜4.2, see ˜J.1.
PSGD (NSL)
Corollary 4.8.
Under ˜3 and 3, Algorithm˜1 with PSGD (NSL) solves Equation˜1.2 exactly in finitely many iterations.
When the learning rate is SRSL, one can explicitly construct the function $Q_{L,\theta^{1}}$ (see ˜J.2); hence, one can bound the number of iterations required to solve the DDIOP associated with an MILP as follows:
Corollary 4.9.
Under ˜3 and 3, Algorithm˜1 with PSGD (SRSL) solves Equation˜1.2 exactly within at most
$\max\left(1,\left(\frac{L(\ell_{\mathrm{sub}})(\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2})}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right)$ iterations.
For the construction of $Q_{L,\theta^{1}}$ for NSL (including SRSL) and the proof of ˜4.2, see ˜J.2.
4.3. Extension: finite-time attainment of the PLF minimum
In this section, we show that, by using PSGD, one can drive the PLF to 0 within finitely many iterations. For a nonempty subset $\Theta^{\prime}\subset\Theta$ and a point $\theta\in\Theta$, define $d(\theta,\Theta^{\prime}):=\inf_{\theta^{\prime}\in\Theta^{\prime}}\|\theta-\theta^{\prime}\|$.
As in the case of the DDIOP for MILPs, by employing PSGD (NSS, NSL), one can attain the minimum value 0 of the PLF in finitely many iterations (Theorems˜N.5 and N.7). In particular, for PSGD (SRSS, SRSL), we can bound the distance between the iterate sequence $\{\theta^{t}\}_{t}$ and the set of PLF minimizers $\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{plf}}$, as well as the number of iterations required to attain the PLF minimum (Theorems˜4.10 and 4.11).
Theorem 4.10.
Suppose that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence obtained by implementing the update rule $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 via PSGD (SRSS). Then: (A) For $T\in\mathbb{Z}_{\geq 1}$,
$d(\theta^{T+1},\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{plf}})^{2}$
$\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right)$
or $\ell_{\mathrm{plf}}(\theta^{T})=0$. (B) The condition $\ell_{\mathrm{plf}}(\theta^{T})=0$ is attained within at most
$\max\left(\frac{4(1+\log 2)^{2}L(\ell_{\mathrm{sub}})^{4}\beta^{2}}{\gamma(\ell_{\mathrm{sub}})^{2}}-2,\left(1+\frac{\mathrm{diam}(\Theta)^{2}}{(2-\sqrt{2})\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}\right)$ iterations.
Theorem 4.11.
Suppose that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence obtained by implementing the update rule $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 via PSGD (SRSL). Then: (A) For $T\in\mathbb{Z}_{\geq 1}$,
$d(\theta^{T+1},\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{plf}})^{2}\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\beta\frac{2(1+\log 2)}{\sqrt{T+2}}\right)$
or $\ell_{\mathrm{plf}}(\theta^{T})=0$. (B) The condition $\ell_{\mathrm{plf}}(\theta^{T})=0$ holds within at most
$\max\left(\frac{4(1+\log 2)^{2}L(\ell_{\mathrm{sub}})^{2}\beta^{2}}{\gamma(\ell_{\mathrm{sub}})^{2}}-2,\left(1+\frac{\mathrm{diam}(\Theta)^{2}L(\ell_{\mathrm{sub}})}{(2-\sqrt{2})\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}\right)$ iterations.
For the proofs of Theorems˜4.10 and 4.11, see Appendix˜N.
5. Proof Overview / Intuition
We provide an intuition for the claim that “for convex piecewise-linear functions whose minimizer set has an interior point, first-order methods such as projected subgradient descent reach the minimum value in finitely many iterations.”
In this paper, we take $\Theta$ to be a convex compact set and consider first-order methods (methods whose iterates are updated solely based on the past iterates and (sub)gradients; e.g., PSGD). In this setting, for a convex piecewise-linear function $\ell:\Theta\to\mathbb{R}$, if the minimizer set $\operatorname*{arg\,min}_{\Theta}\ell$ has a relative interior point, then one can construct another convex piecewise-linear function $\widetilde{\ell}$ by decreasing the function values only in the interior of $\operatorname*{arg\,min}_{\Theta}\ell$. Moreover, since one may construct $\widetilde{\ell}$ so that $\ell=\widetilde{\ell}$ holds outside $\operatorname*{arg\,min}_{\Theta}\ell$, the iterates produced by a first-order method applied to $\ell$ and those produced when applied to $\widetilde{\ell}$ coincide (and so do the subgradients) until the iterates reach $\operatorname*{arg\,min}_{\Theta}\ell$.
[…]
Figure 2. Schematic illustration of reaching the minimum value $\min_{\Theta}\ell=0$ in finitely many iterations when applying the (sub)gradient method to the convex piecewise-linear function $\ell(\theta)=\max(|\theta|-1,0)$.
We illustrate this behavior via a concrete example. Consider $\ell(\theta)=\max(|\theta|-1,0)$. Then $\operatorname*{arg\,min}\ell=[-1,1]$, and the minimizer set has an interior point. Let $\widetilde{\ell}(\theta)=|\theta|-1$. Since $\ell(\theta)=\widetilde{\ell}(\theta)$ holds for $|\theta|>1$, the iterates of any first-order method are identical for the two objectives until they enter [1,1][-1,1]. Furthermore, since $\min_{\Theta}\ell=0$ and $\min_{\Theta}\widetilde{\ell}<0$, we can define $\gamma:=\min_{\Theta}\ell-\min_{\Theta}\widetilde{\ell}>0$ (the gap is depicted as $\gamma$ in Figure˜2).
On the other hand, for first-order methods such as projected subgradient descent, standard convergence guarantees imply that $\widetilde{\ell}(\theta^{t})$ converges to $\min_{\Theta}\widetilde{\ell}$.” Hence, there exists a finite iteration index $T$ such that $\widetilde{\ell}(\theta^{T})\leq\min_{\Theta}\widetilde{\ell}+\gamma=\min_{\Theta}\ell$. Since the iterates coincide until the minimum is attained, the same $T$ yields $\ell(\theta^{T})=\min_{\Theta}\ell$, i.e., the minimum value of $\ell$ is attained in finitely many iterations. Finally, the main object of interest in this paper, the suboptimality loss $\ell_{\mathrm{sub}}$, satisfies the following proposition:
Proposition 5.1.
Assume ˜3 and 3. Then: (A) it is piecewise-linear; (B) the minimum value of the suboptimality loss is 0; and (C) the set of minimizers has an interior point.
See ˜L.3 for the proof. By ˜3 and 5 and ˜3, the suboptimality loss $\ell_{\mathrm{sub}}$ is convex, Lipschitz, and piecewise-linear; moreover, when $\Theta=\Delta^{d-1}$, its minimizer set has an interior point (for Lebesgue-almost every true weight $\theta^{*}$). Therefore, by an argument analogous to the above intuition, one expects that first-order methods including PSGD attain $\ell_{\mathrm{sub}}=0$ in finitely many iterations (i.e., solve the DDIOP for MILPs exactly), as formalized in Theorem˜4.3.
6. Numerical Experiment
In this section, under ˜3, we report numerical results comparing Algorithm˜1 with existing methods. Implementation details and the computational device used are provided in ˜O.1.
When the objective function is linear and $X(s)$ is a convex region, it is known that one can solve Chan et al. (2023, (15)) by taking the number of point clouds $T$ uniformly. We refer to this method as CHAN.
Table 2. Computational time per iteration, or equivalently, the computational cost required for a single call to the forward problem, for each method when $\Theta=\Delta^{d-1}$. The constant $\mathrm{Opt}$ denotes the computational complexity of running Equation˜$1.$
Method Computational complexity
UPA $N\mathrm{Opt}$
RPA $N\mathrm{Opt}$
CHAN at least $N\mathrm{Opt}$
PSGD $N\mathrm{Opt}+O(d\log d)$
When $\Theta=\Delta^{d-1}$, the computational time per iteration, or equivalently, the computational cost required for a single call to the forward problem, is summarized in Table˜2. The computational complexity of one iteration of Algorithm˜1 is the sum of the complexity of Equation˜1.1 and the complexity of the projection $\operatorname{Proj}_{\Delta^{d-1}}$, namely $O(d\log d)$ (Wang and Carreira-Perpinán, 2013). Since the present experiments have dimension $d\leq 8$, we may regard the runtime of methods other than CHAN as comparable to, or faster than, that of CHAN.
6.1. Linear programming
Figure 3. Worst-case behavior of $\ell_{\mathrm{plf}}(\theta^{t}(\mathcal{D}))+0.1$ for the LP setting.
The LP setup is described in detail in ˜O.2. Under this setup, we compare PSGD (SRSL), PSGD (Polyak), UPA, RPA, and CHAN.
Each experiment consists of the following five steps. (Step 1) Sample $\theta^{*}\in\Delta^{d-1}$ according to the uniform distribution on $\Delta^{d-1}$. (Step 2) Sample $s^{(n)}\in\mathcal{S}$ according to the random variable specified in ˜O.2. (Step 3) For $n=1,\ldots,N$, set $a^{(n)}=a^{*}(\theta^{*},s^{(n)})$. (Step 4) Run PSGD (SRSL), PSGD (Polyak), UPA, RPA, and CHAN with the number of iterations (or the number of acquired data points) set to $500$. (Step 5) Repeat (Step 1)–(Step 4) $100$ times.
Remark 6.1.
For UPA and CHAN, when these steps are executed only at points indexed by an integer $j$ (for $t<500$) and represented by $|\Theta_{j}|$, we linearly interpolated the loss values between such points.
We summarize the results as follows. Figure˜3 shows, for $d=4,6,8$, the worst-case trajectories of $\ell_{\mathrm{plf}}(\theta^{t}(\mathcal{D}))+0.1$. For the PLF minimization problem, Table˜1 suggests that PSGD (SRSL) is the most effective method for reducing the PLF, achieving PLF minimization in fewer than $1/7$ of the iterations required by UPA, RPA, and CHAN. In particular, when $d\geq 6$, once training with PSGD (SRSL) is complete, PSGD (SRSL) improves the PLF by more than two orders of magnitude compared with the existing algorithms.
By Table˜1, as the dimension $d$ increases, UPA, RPA, and CHAN are expected to move toward the upper right. In contrast, we observe that PSGD (SRSL) is less sensitive to the dimension than the other methods.
Moreover, PSGD (SRSL) drives the PLF to 0 with probability $100\%$. This phenomenon can be theoretically justified by Theorem˜M.3.
Remark 6.2.
The behavior of the suboptimality loss is described in detail in Figure˜6. From a theoretical viewpoint, PSGD (SRSS, SRSL) is the most effective method for reducing the suboptimality loss (see ˜N.3). However, in our experiments, we found that both PSGD (SRSL) and PSGD (Polyak) are efficient.
It remains necessary to investigate why PSGD (Polyak) fails to attain the minimum value of the PLF. Under PSGD (Polyak), the suboptimality loss value becomes extremely small, and consequently the update step size also becomes extremely small. As a result, the update magnitude becomes negligible, preventing the method from attaining the PLF minimum value 0.
6.2. Single-machine weighted sum of completion times scheduling
Figure 4. Worst-case behavior of $\ell_{\mathrm{plf}}(\theta^{t}(\mathcal{D}))+0.1$ for the single-machine scheduling instance.
The single-machine weighted sum of completion times scheduling problem $1|r_{j}|\sum\theta_{j}C_{j}$ is an ILP. The experimental setup for $1|r_{j}|\sum\theta_{j}C_{j}$ is given in ˜O.3.
We conducted experiments following the same steps (Step 1)–(Step 5) as in ˜6.1, with $t\leq 1000$ and excluding CHAN. We summarize the results as follows. Figure˜4 reports, for $d=4,6,8$, the worst-case trajectories of $\ell_{\mathrm{plf}}(\theta^{t}(\mathcal{D}))+0.1$. As in ˜6.1, we confirm that PSGD (SRSL) exhibits weaker dependence on the dimension than existing methods and minimizes the PLF in fewer than $1/10$ of the iterations required by the other methods. The runtimes of each method are reported in Tables˜9, 9 and 9. PSGD (SRSL) terminates within $2.27$ seconds for $d=4,6,8$.
Remark 6.3.
The behavior of the suboptimality loss is shown in Figure˜5.
7. Conclusion
To solve the DDIOP for MILPs, we focused on the suboptimality loss, which is Lipschitz continuous and convex. By leveraging its convex piecewise-linear structure and the existence of an interior point in its minimizer set, we showed that a broad class of gradient-based optimization methods, including PSGD, attains the minimum value 0 of the suboptimality loss in finitely many iterations, and hence solves the DDIOP for MILPs exactly. As a further consequence, we proved that PSGD attains the minimum value of the PLF in finitely many iterations. We also derived an explicit upper bound on the number of iterations required for PSGD to reach the minimum in finite time, and verified the finite-time attainment behavior through numerical experiments.
Several directions for future work are as follows. First, it is of interest to develop algorithms that can efficiently learn inverse optimization problems when the given data contain noise. Real-world data tend to be noisy (cf. Aswani et al., 2018). Establishing methods that simultaneously achieve computational efficiency and robustness to noise is expected to broaden the applicability to practical optimization problems. Second, the iteration-complexity upper bound for finite-time attainment obtained in this paper is expressed via the constant $\gamma(\ell_{\mathrm{sub}})$. The quantity $\gamma(\ell_{\mathrm{sub}})$ depends on the observed data $\mathcal{D}$ and on the structure of the forward problem, and is generally difficult to evaluate explicitly. In particular, when the diversity of states increases or the sample size $N$ becomes large, $\gamma(\ell_{\mathrm{sub}})$ may become small, which can lead to a large upper bound on the required number of iterations. Obtaining lower bounds on $\gamma(\ell_{\mathrm{sub}})$ and deriving sharper complexity estimates based on an analysis of its statistical behavior with respect to $N$ remain important topics for future research.
Appendix A Definitions of UPA and RPA
We briefly review UPA and RPA. For a subset $\Theta^{\prime}\subset\Delta^{d-1}$, define
$\psi(\Theta^{\prime}):=\operatorname*{arg\,min}_{\Theta^{\prime}}\ell_{\mathrm{plf}}\,.$
In UPA, we define
$\Theta_{k}:=\left\{\left(\frac{2k_{1}+1}{2k+d},\ldots,\frac{2k_{d}+1}{2k+d}\right)\,\middle|\,\forall i,\ k_{i}\in\mathbb{Z}_{\geq 0},\ \sum_{i}k_{i}=k\right\}\,.$
UPA returns $\psi(\Theta_{k})$. In RPA, letting $\Theta_{\mathrm{rand}}$ be a set of $k$ random points sampled from $\Delta^{d-1}$, RPA returns $\psi(\Theta_{\mathrm{rand}})$.
Appendix B Related works
B.1. Methods for smoothing the prediction loss of optimal solutions
The input-perturbation approach of Berthet et al. (2020) requires multiple steps in order to obtain a solution and guarantee its uniqueness. Moreover, since each gradient computation involves a Monte Carlo procedure (suppose that $M$ samples are drawn), the overall computational cost of learning becomes substantial. This effectively means that one must solve $M$ times more linear programs. In addition, the required number of samples $M$ for gradient estimation must itself be selected or estimated.
For LPs, Wilder et al. (2019) proposed a method that adds a regularization term to the objective function. However, to apply this method, $X(s)$ must be described by linear constraints. Hence, it is not directly applicable when $X(s)$ is nonconvex, as in ILPs or MILPs. To address this limitation, Ferber et al. (2020) proposed reducing an MILP to an LP by using Gomory cuts (Gomory, 1960) and globally valid cuts (Balas et al., 1996). Smoothing the PLF within the framework of Ferber et al. (2020) provides one approach to solving the PLF minimization problem via backpropagation (cf. (Ferber et al., 2023)). However, to the best of our knowledge, the method proposed in Ferber et al. (2023) does not provide an approximation-error analysis, nor does it establish theory guaranteeing uniqueness of the solution. Furthermore, it remains unclear what level of accuracy and computational effort is required for the Gomory cuts and globally valid cuts used in Ferber et al. (2020).
B.2. The case where the objective is strongly convex and $L$-smooth and the constraint set is convex
Assume that the objective function in Equation˜1.1 is $\mu$-strongly convex and $L$-smooth, and that $X(s)$ is convex. In this case, minimizing the PLF is equivalent to minimizing the suboptimality loss (see Appendix˜C). Moreover, by running projected gradient descent on the suboptimality loss, the resulting sequence of PLF values converges linearly (see Appendix˜D).
B.3. PLF minimization via the suboptimality loss
Assume that the feature vector $a^{*}(\theta,s^{(n)})$ returns, among the solutions to
$\operatorname*{arg\,max}_{f(x)\in f(X(s))}\ \theta^{\top}f(x),$
the one that is minimal in the lexicographic order. Under this assumption, Kitaoka and Eto (2023b) showed that $a^{*}(\theta,s^{(n)})=a^{(n)}$ holds whenever a subgradient of the suboptimality loss is equal to 0.
B.4. SPO loss
In the context of inverse optimization, there is a line of work using the SPO loss Mohajerin Esfahani et al. (2018); Elmachtoub and Grigas (2022). The constant $\gamma(\ell_{\mathrm{sub}})$ can be upper-bounded by employing the SPO loss Mohajerin Esfahani et al. (2018); Elmachtoub and Grigas (2022). For details, see ˜N.4.
Appendix C Prediction loss and suboptimality loss: the case of strongly convex and smooth objectives with convex constraints
In this section, we establish the following result.
Proposition C.1.
Let the weight space $\Theta$ be a convex set. Assume that, for any $\theta\in\Theta$, the function $\theta^{\top}f(x)$ is $\mu$-strongly convex and differentiable in $x$, and that, for every $n$, there exists $x^{(n)}\in X(s^{(n)})$ satisfying $a^{(n)}=f(x^{(n)})$. Then, minimizing the suboptimality loss is equivalent to minimizing the PLF; namely,
$\ell_{\mathrm{sub}}(\theta)=0\Leftrightarrow\ell_{\mathrm{plf}}(\theta)=0\Leftrightarrow\forall n,\ a^{(n)}=a^{*}(\theta,s^{(n)})\,.$
Before proving the above, we recall the following known proposition.
Proposition C.2.
Mohajerin Esfahani et al. (2018, Proposition 2.5) Let the weight space $\Theta$ be a convex set. Assume that, for any $\theta\in\Theta$, the function $\theta^{\top}f(x)$ is $\mu$-strongly convex and differentiable in $x$, and that, for every $n$, there exists $x^{(n)}\in X(s^{(n)})$ satisfying $a^{(n)}=f(x^{(n)})$. Then,
$\ell_{\mathrm{sub}}(\theta)\geq\frac{\mu}{2N}\sum_{n=1}^{N}\left\|x^{*}(\theta,s^{(n)})-x^{(n)}\right\|^{2}$
holds.
Proof of Appendix˜C.
Assuming that the mapping $f$ is $L(f)$-Lipschitz continuous, we obtain
$L(f)^{2}\ell_{\mathrm{sub}}(\theta)$ $\geq L(f)^{2}\frac{\mu}{2N}\sum_{n=1}^{N}\left\|x^{*}(\theta,s^{(n)})-x^{(n)}\right\|^{2}$
$\geq\frac{\mu}{2N}\sum_{n=1}^{N}\left\|a^{*}(\theta,s^{(n)})-a^{(n)}\right\|^{2}=\frac{\mu}{2}\ell_{\mathrm{plf}}(\theta)$
On the other hand, by the Cauchy–Schwarz inequality,
$\ell_{\mathrm{sub}}(\theta)\leq\sup_{\theta\in\Theta}\|\theta\|\,\ell_{\mathrm{plf}}(\theta)^{\nicefrac{{1}}{{2}}}\,.$
Combining the above, we conclude that
$\ell_{\mathrm{sub}}(\theta)=0\Leftrightarrow\ell_{\mathrm{plf}}(\theta)=0\Leftrightarrow\forall n,\ a^{(n)}=a^{*}(\theta,s^{(n)})$
holds. ∎
Appendix D Projected gradient descent when the objective is strongly convex and smooth and the constraints are convex
In this section, we establish the following result.
Proposition D.1.
Let the weight space $\Theta$ be a convex set. Assume that, for any $\theta\in\Theta$, the function $\theta^{\top}f(x)$ is $\mu$-strongly convex and differentiable in $x$, and that, for every $n$, there exists $x^{(n)}\in X(s^{(n)})$ satisfying $a^{(n)}=f(x^{(n)})$. Then, if one runs projected gradient descent on the iterate sequence $\{\theta^{t}\}$ with step size $\alpha_{t}=1/\beta$, the PLF converges linearly; namely,
$\ell_{\mathrm{plf}}(\theta^{t})\leq\frac{2\ell_{\mathrm{sub}}(\theta^{1})}{\mu}\exp\left(-\frac{\mu}{4\beta}(t-1)\right)\,.$
proof.
By Hazan (2019, Theorem 2.6), $\ell_{\mathrm{sub}}(\theta^{t})$ admits the bound
$\ell_{\mathrm{sub}}(\theta^{t})\leq\ell_{\mathrm{sub}}(\theta^{1})\exp\left(-\frac{\mu}{4\beta}(t-1)\right)\,.$
By Appendix˜C, it follows that
$\ell_{\mathrm{plf}}(\theta^{t})\leq\frac{2}{\mu}\ell_{\mathrm{sub}}(\theta^{t})\leq\frac{2\ell_{\mathrm{sub}}(\theta^{1})}{\mu}\exp\left(-\frac{\mu}{4\beta}(t-1)\right)\,.$
Appendix E Regret analysis for the suboptimality loss
Let $\Theta$ be a nonempty subset of $\mathbb{R}^{d}$. Let $\ell\colon\Theta\to\mathbb{R}$ be a loss function. Consider a sequence $\{\theta^{t}\}_{t=1}^{T}\subset\Theta$. We define the regret by
$\mathrm{Regret}(T)=\sum_{t=1}^{T}\left(\ell(\theta^{t})-\min_{\theta\in\Theta}\ell(\theta)\right)\,.$
In what follows, we review existing work on regret analyses for the suboptimality loss. The results are summarized in Table˜3.
Table 3. Performance comparison of methods for solving the DDIOP for MILPs (Equation˜1.2). The integer $T$ denotes the number of iterations or the number of optimization calls. The integer $t$ denotes the iteration index or the optimization-call index.
Method Regret for the suboptimality loss
MWU $O\left(L(\ell_{\mathrm{sub}})\log d\sqrt{T}\right)$, provided that $\Theta=\Delta^{d-1}$
(Arora et al., 2012) (Bärmann et al., 2018, Theorem 3.5)
PSGD (online gradient descent) $O\left(\mathrm{diam}(\Theta)L(\ell_{\mathrm{sub}})\log d\sqrt{T}\right)$
(with step size (Bärmann et al., 2018, Theorem 3.11)
$\mathrm{diam}(\Theta)L(\ell_{\mathrm{sub}})^{-1}t^{-1/2}$)
Besbes et al. (2021, 2025) $O\left(d^{4}\log T\right)$, provided that $\Theta$ is the unit sphere and $L(\ell_{\mathrm{sub}})\leq 1$
Gollapudi et al. (2021) $O\left(d\log T\right)$, provided that $\Theta$ is the unit ball and $L(\ell_{\mathrm{sub}})\leq 1$
ONS $O\left(\mathrm{diam}(\Theta)L(\ell_{\mathrm{sub}})d\log(T/d)\right)$
(Hazan et al., 2007) (Sakaue et al., 2025, Theorem 3.1)
MetaGrad $O\left(\mathrm{diam}(\Theta)L(\ell_{\mathrm{sub}})d\log(T/d)\right)$
(van Erven and Koolen, 2016; (Sakaue et al., 2025, Theorem 4.1)
van Erven et al., 2021)
Cf. lower bound $\Omega(d)$ (Sakaue et al., 2025, §5)
Appendix F Online and offline optimization
Let $\Theta$ be a nonempty subset of $\mathbb{R}^{d}$. Let $\ell\colon\Theta\to\mathbb{R}$ be a function. Consider a sequence $\{\theta^{t}\}_{t=1}^{T}\subset\Theta$. Then, the following holds:
$\min_{t=1,\ldots,T}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)\leq\frac{1}{T}\sum_{t=1}^{T}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)\leq\frac{\mathrm{Regret}(T)}{T}\,.$ (F.1)
By applying the existing regret bounds reviewed in Appendix˜E to Equation˜F.1, one can upper-bound the best-iterate performance for the suboptimality loss. For the resulting bounds, see Table˜4.
Appendix G Comparison tables
G.1. DDIOP for MILPs
A performance comparison of methods for solving the DDIOP for MILPs (Equation˜1.2) is provided in Table˜4.
Table 4. Performance comparison of methods for solving the DDIOP for MILPs (Equation˜1.2). The integer $T$ denotes the number of iterations or the number of optimization calls. The integer $t$ denotes the iteration index or the optimization-call index.
Method Suboptimality loss
UPA $O\!\left(\mathrm{diam}(\Theta)\,L(\ell_{\mathrm{sub}})\,T^{-\nicefrac{{1}}{{(d-1)}}}\right)$
(cf. Flatto and Newman, 1977, Theorem 2.1)
RPA $O_{\mathbb{P}}\!\left(\mathrm{diam}(\Theta)\,L(\ell_{\mathrm{sub}})\,\left(\log T/T\right)^{\nicefrac{{1}}{{(d-1)}}}\right)$
(cf. Reznikov and Saff, 2015, Corollary 2.3)
MWU $O\left(L(\ell_{\mathrm{sub}})\log d/\sqrt{T}\right)$, provided that $\Theta=\Delta^{d-1}$
(Arora et al., 2012) (Bärmann et al., 2018, Theorem 3.5)
Besbes et al. (2021, 2025) $O\left(d^{4}\log T/T\right)$, provided that $\Theta$ is the unit sphere and $L(\ell_{\mathrm{sub}})\leq 1$
Gollapudi et al. (2021) $O\left(d\log T/T\right)$, provided that $\Theta$ is the unit ball and $L(\ell_{\mathrm{sub}})\leq 1$
ONS $O\left(\mathrm{diam}(\Theta)\,L(\ell_{\mathrm{sub}})\,d\log(T/d)/T\right)$
(Hazan et al., 2007) (Sakaue et al., 2025, Theorem 3.1)
MetaGrad $O\left(\mathrm{diam}(\Theta)\,L(\ell_{\mathrm{sub}})\,d\log(T/d)/T\right)$
(van Erven and Koolen, 2016; (Sakaue et al., 2025, Theorem 4.1)
van Erven et al., 2021)
PSGD $O\left(\mathrm{diam}(\Theta)\,L(\ell_{\mathrm{sub}})\,\log d/\sqrt{T}\right)$
(with step size (Bärmann et al., 2018, Theorem 3.11)
$\mathrm{diam}(\Theta)L(\ell_{\mathrm{sub}})^{-1}t^{-1/2}$)
PSGD 0 if $T\geq\max\left(1,\left(\frac{\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}L(\ell_{\mathrm{sub}})^{2}}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right)$
(with step size $\beta t^{-1/2}$) (Ours, ˜4.2 and J.1)
PSGD 0 if $T\geq\max\left(1,\left(\frac{L(\ell_{\mathrm{sub}})\,(\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2})}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right)$
(with step length $\beta t^{-1/2}$) (Ours, ˜4.2 and J.2)
G.2. PLF minimization
A performance comparison of methods for solving the minimum value of the PLF for MILPs is provided in Table˜5.
Table 5. Performance comparison of methods for solving the DDIOP for MILPs (Equation˜1.2). The notation is the same as in Table˜4.
Method Distance between the set where the PLF equals 0 and the learned weight $\theta^{T}$ (when the PLF is nonzero)
UPA $O\!\left(\mathrm{diam}(\Theta)\,T^{-\nicefrac{{1}}{{(d-1)}}}\right)$
RPA $O_{\mathbb{P}}\!\left(\mathrm{diam}(\Theta)\,\left(\log T/T\right)^{\nicefrac{{1}}{{(d-1)}}}\right)$
PSGD $\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T-1}-1\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+1}}\right)$,
(with step size $\beta t^{-1/2}$) Theorem˜4.10
PSGD $\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T-1}-1\right)\left(2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\beta\frac{2(1+\log 2)}{\sqrt{T+1}}\right)$,
(with step length $\beta t^{-1/2}$) Theorem˜4.11
Appendix H Additional remarks on Assumption 3
Remark H.1.
We explain why it is preferable to exclude 0 from the weight space $\Theta$. At the origin 0, one has $\ell_{\mathrm{sub}}(0)=0$, and thus 0 attains the minimum value of the suboptimality loss. Consequently, for the DDIOP for MILPs (Equation˜1.2), $\theta=0$ is a solution.
However, the optimizer $x^{*}(0,s^{(n)})$ may be any point in $X(s^{(n)})$; hence, unless $X(s^{(n)})$ is a singleton, the optimizer is not uniquely determined. Therefore, even if the true weight were $\theta^{*}=0$, Assumption 3 would not be satisfied. That is, even if learning returns $\theta=0$, such a $\theta$ does not coincide with the true weight $\theta^{*}$. This issue is common to many IOPs for LPs, regardless of whether the setting is online or offline Bertsimas et al. (2015); Mohajerin Esfahani et al. (2018); Bärmann et al. (2018); Chen and Kılınç-Karzan (2020); Sun et al. (2023); Kitaoka and Eto (2023a, b).
Appendix I Proof of Lemma 3
We introduce the following notation:
$\Psi$ $:=\bigcap_{n=1}^{N}\bigcup_{\xi^{(n)}\in Y^{(n)}}\Theta(\xi^{(n)},Y^{(n)})\,$,
$\Theta(\xi^{(n)},Y^{(n)})$ $:=\left\{\theta\in\Theta\,\middle|\,\forall b\in Y^{(n)}\setminus\{\xi^{(n)}\},\ \theta^{\top}\xi^{(n)}>\theta^{\top}b\right\}\,.$
For any $n=1,\ldots,N$ and any $\xi^{(n)}\in Y^{(n)}$, the set $\Theta(\xi^{(n)},Y^{(n)})$ is open in the relative topology on $\Theta$. That is, for any $\theta^{*}\in\Psi$, there exist $\varepsilon(\theta^{*})>0$ and points $\xi^{(n)}\in Y^{(n)}$ (for each $n$) such that
$\left\{\theta\in\Delta^{d-1}\,\middle|\,\|\theta-\theta^{*}\|<\varepsilon(\theta^{*})\right\}\subset\bigcap_{n=1}^{N}\Theta(\xi^{(n)},Y^{(n)})$ (I.1)
holds.
Next, for $D\subset Y^{(n)}$, define
$\Theta(D,Y^{(n)}):=\left\{\theta\in\Delta^{d-1}\,\middle|\,\forall a_{1},a_{2}\in D,\ \forall a_{3}\in Y^{(n)}\setminus D,\ \theta^{\top}a_{1}=\theta^{\top}a_{2}>\theta^{\top}a_{3}\right\}.$
Then,
$\Delta^{d-1}=\bigcup_{D\subset Y^{(n)}}\Theta(D,Y^{(n)}).$
We further define
$\mathcal{M}(Y^{(n)})$ $:=\left\{D\subset Y^{(n)}\,\middle|\,\dim\Theta(D)=d-1\right\}\,.$
Lemma I.1.
For any $D\in\mathcal{M}(Y^{(n)})$, the set $D$ is a singleton.
proof.
Assume that $D$ contains two distinct points, and take $a_{1},a_{2}\in D$. If there exists $c\neq 0$ such that $a_{1}-a_{2}=(c,\ldots,c)$, then, since $\theta\in\Delta^{d-1}$,
$0=\theta^{\top}(a_{1}-a_{2})=\theta^{\top}(c,\ldots,c)=c$
holds, which contradicts $c\neq 0$. Hence, we may assume that $a_{1}-a_{2}$ and $(1,\ldots,1)$ are linearly independent. However,
$\Theta(D,Y^{(n)})\subset\left\{\theta\in\mathbb{R}^{d}\,\middle|\,\theta^{\top}(a_{1}-a_{2})=0,\ \theta^{\top}(1,\ldots,1)=1\right\},$
implying that $\dim\Theta(D,Y^{(n)})\leq d-2$. This contradicts $D\in\mathcal{M}(Y^{(n)})$. ∎
Lemma I.2.
Let $D_{1},D_{2}\subset Y^{(n)}$. If $D_{1}\neq D_{2}$, then $\Theta(D_{1},Y^{(n)})\cap\Theta(D_{2},Y^{(n)})=\emptyset$.
proof.
Suppose that $D_{1}\neq D_{2}$ and that there exists $\theta\in\Theta(D_{1},Y^{(n)})\cap\Theta(D_{2},Y^{(n)})$. By symmetry, we may assume $D_{1}\setminus D_{2}\neq\emptyset$. Take $a\in D_{1}\setminus D_{2}$ and $b\in D_{2}$. Since $\theta\in\Theta(D_{2},Y^{(n)})$, we have $\theta^{\top}b>\theta^{\top}a$. On the other hand, since $\theta\in\Theta(D_{1},Y^{(n)})$, we have $\theta^{\top}b\leq\theta^{\top}a$. This is a contradiction. ∎
Proof of ˜3.
By Appendix˜I, we have
$\Delta^{d-1}\setminus\bigcup_{D\in\mathcal{M}(Y^{(n)})}\Theta(D,Y^{(n)})=\bigcup_{\begin{subarray}{c}D\subset Y^{(n)}\\ \dim\Theta(D,Y^{(n)})<d-1\end{subarray}}\Theta(D,Y^{(n)}).$
If $\dim\Theta(D,Y^{(n)})<d-1$, then the convex set $\Theta(D,Y^{(n)})$ has (Lebesgue) measure zero in $\mathbb{R}^{d}$ (equivalently, zero (d1)(d-1)-dimensional measure relative to the simplex). Since the power set of $Y^{(n)}$ is finite, it follows that
$\Delta^{d-1}\setminus\bigcup_{D\in\mathcal{M}(Y^{(n)})}\Theta(D,Y^{(n)})$
has measure zero. Moreover,
$\Delta^{d-1}\setminus\Psi=\bigcup_{n=1,\ldots,N}\left(\Delta^{d-1}\setminus\bigcup_{D\in\mathcal{M}(Y^{(n)})}\Theta(D,Y^{(n)})\right),$
and hence $\Delta^{d-1}\setminus\Psi$ has measure zero. This proves the lemma. ∎
Appendix J Finite-time exact solvability of the DDIOP for MILPs
J.1. PSGD (NSS)
For PSGD (NSS), the following convergence rate is well known.
Proposition J.1 (e.g. Beck, 2017, Theorem 8.25).
Let $\Theta$ be a bounded, closed, and convex set, and let $L>0$, $\theta^{1}\in\Theta$, and $T\in\mathbb{Z}_{\geq 1}$. Then, (A) for any $L$-Lipschitz convex function $\ell\colon\Theta\to\mathbb{R}$, the sequence $\{\theta^{t}\}_{t=1}^{T}$ generated by PSGD (NSS) satisfies
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\leq\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\sum_{t=1}^{T}\beta_{t}}+\frac{L^{2}}{2}\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\,.$
In particular, (B) there exists a function $Q_{L,\theta^{1}}\colon\mathbb{R}_{>0}\to\mathbb{R}$ satisfying ˜4.1.
Proposition J.2.
Assume that ˜3 and 3 hold. Let $0<q<1$ and $\beta>0$, and suppose that the step size is given by $\beta_{t}=\beta t^{-q}$. Define
$Q_{L,\theta^{1}}(\varepsilon)=\begin{cases}\left(\frac{\mathrm{diam}(\Theta)^{2}+L^{2}\beta^{2}(1-2q)}{2\beta(1-q)\varepsilon}\right)^{1/q},&0<q<1/2,\\ \left(\frac{\mathrm{diam}(\Theta)^{2}+L^{2}\beta^{2}(1-2q)}{2\beta(1-q)\varepsilon}\right)^{1/(1-q)},&1/2<q<1,\\ \max\left(1,\left(\frac{\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}L^{2}}{\beta\varepsilon}\right)^{2}-2\right),&q=1/2\,.\end{cases}$ (J.1)
Then $Q_{L,\theta^{1}}$ satisfies ˜4.1.
By Theorem˜4.3 and ˜J.1, we obtain the following corollary.
Corollary J.3.
Assume that ˜3 and 3 hold. Then, if we implement Algorithm˜1 using PSGD in the present setting, the DDIOP Equation˜1.2 is solved exactly within at most
$\begin{cases}\left(\frac{\mathrm{diam}(\Theta)^{2}+L(\ell_{\mathrm{sub}})^{2}\beta^{2}(1-2q)}{2\beta(1-q)\gamma(\ell_{\mathrm{sub}})}\right)^{1/q},&0<q<1/2,\\ \left(\frac{\mathrm{diam}(\Theta)^{2}+L(\ell_{\mathrm{sub}})^{2}\beta^{2}(1-2q)}{2\beta(1-q)\gamma(\ell_{\mathrm{sub}})}\right)^{1/(1-q)},&1/2<q<1,\\ \max\left(1,\left(\frac{\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}L(\ell_{\mathrm{sub}})^{2}}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right),&q=1/2\end{cases}$
iterations.
We next prepare the proof of ˜J.1.
Proposition J.4 (Beck, 2017, Lemmas 8.11 and 8.24).
Let $\Theta$ be a bounded, closed, and convex set, and let $\theta^{1}\in\Theta$ and $T\in\mathbb{Z}_{\geq 1}$. Let $\{\alpha_{t}\}_{t=1}^{T}$ be a step-size sequence. Let $\ell\colon\Theta\to\mathbb{R}$ be a Lipschitz continuous convex function. Take $\theta^{*}\in\operatorname*{arg\,min}_{\Theta}\ell$. Let $\{\theta^{t}\}_{t=1}^{T+1}$ be the sequence produced by PSGD. Then, for any $t=1,\ldots,T$,
$\|\theta^{t+1}-\theta^{*}\|^{2}$ $\leq\|\theta^{t}-\theta^{*}\|^{2}-2\alpha_{t}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)+\alpha_{t}^{2}\|\nabla\ell(\theta^{t})\|^{2},$ (J.2)
$\sum_{t=1}^{T}\alpha_{t}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)$ $\leq\frac{1}{2}\|\theta^{1}-\theta^{*}\|^{2}+\frac{1}{2}\sum_{t=1}^{T}\alpha_{t}^{2}\|\nabla\ell(\theta^{t})\|^{2}\,.$ (J.3)
Proposition J.5.
Let $\delta>0$. Then, for any $T\in\mathbb{Z}_{\geq 1}$,
$\frac{\delta+\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\leq\frac{2(\delta+1+\log 2)}{\sqrt{T+2}}\,.$
proof.
For $T=1$, the inequality holds.
Let $T\in\mathbb{Z}_{\geq 2}$ and set $k:=\lceil T/2\rceil$. Then,
$\sum_{t=k}^{T}t^{-1}$ $\leq 1+\int_{k}^{T}z^{-1}\,dz=1+\log\left(\frac{T}{k}\right)\leq 1+\log\left(\frac{T}{T/2}\right)=1+\log 2\,.$
On the other hand,
$\sum_{t=k}^{T}t^{-1/2}$ $\geq\int_{k}^{T+1}z^{-1/2}\,dz=2\sqrt{T+1}-2\sqrt{k}\geq 2\sqrt{T+1}-2\sqrt{T/2}$
$=2\frac{T+1-T/2}{\sqrt{T+1}+\sqrt{T/2}}\geq\frac{T+2}{2\sqrt{T+2}}=\frac{\sqrt{T+2}}{2}\,.$
Combining these bounds yields the claim. ∎
Proof of ˜J.1.
Assume $0<q<1$ and $q\neq 1/2$. Then,
$\sum_{t=1}^{T}\beta_{t}\geq\beta(1-q)T^{1-q},\quad\sum_{t=1}^{T}\beta_{t}^{2}\leq\beta^{2}(1-2q)T^{1-2q}.$ (J.4)
Hence,
$\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\sum_{t=1}^{T}\beta_{t}}+\frac{L^{2}}{2}\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}$ $\leq\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{1-q}}+\frac{L^{2}}{2}\frac{\beta^{2}(1-2q)T^{1-2q}}{\beta(1-q)T^{1-q}}$
$=\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{1-q}}+\frac{L^{2}\beta}{2}\frac{1-2q}{(1-q)T^{q}}\,.$ (J.5)
If $0<q<1/2$, then by Equation˜J.5 and q<1qq<1-q,
$\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{1-q}}+\frac{L^{2}\beta}{2}\frac{1-2q}{(1-q)T^{q}}$ $\leq\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{q}}+\frac{L^{2}\beta}{2}\frac{1-2q}{(1-q)T^{q}}$
$=\frac{\mathrm{diam}(\Theta)^{2}+L^{2}\beta^{2}(1-2q)}{2\beta(1-q)}T^{-q}\,.$
By ˜J.1,
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\leq\frac{\mathrm{diam}(\Theta)^{2}+L^{2}\beta^{2}(1-2q)}{2\beta(1-q)}T^{-q}.$
Therefore, ˜4.1 holds for $0<q<1/2$.
If $1/2<q<1$, then by Equation˜J.5 and 1q<q1-q<q,
$\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{1-q}}+\frac{L^{2}\beta}{2}\frac{1-2q}{(1-q)T^{q}}$ $\leq\frac{1}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\beta(1-q)T^{1-q}}+\frac{L^{2}\beta}{2}\frac{1-2q}{(1-q)T^{1-q}}$
$=\frac{\mathrm{diam}(\Theta)^{2}+L^{2}\beta^{2}(1-2q)}{2\beta(1-q)}T^{-(1-q)}\,.$
Thus, ˜4.1 holds for $1/2<q<1$.
Finally, consider $q=1/2$. Summing Equation˜J.2 over $t=\lceil T/2\rceil,\lceil T/2\rceil+1,\ldots,T$ yields
$\sum_{t=\lceil T/2\rceil}^{T}2\beta_{t}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)$ $\leq\|\theta^{\lceil T/2\rceil}-\theta^{*}\|^{2}-\|\theta^{T+1}-\theta^{*}\|^{2}+\sum_{t=\lceil T/2\rceil}^{T}\beta_{t}^{2}\|\nabla\ell(\theta^{t})\|^{2}$
$\leq\mathrm{diam}(\Theta)^{2}+\sum_{t=\lceil T/2\rceil}^{T}\frac{\beta^{2}}{t}\|\nabla\ell(\theta^{t})\|^{2}$
$\leq\mathrm{diam}(\Theta)^{2}+\beta^{2}L^{2}\sum_{t=\lceil T/2\rceil}^{T}\frac{1}{t}\,.$
On the other hand,
$\sum_{t=\lceil T/2\rceil}^{T}2\beta_{t}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)\geq 2\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\right).$
Therefore,
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell$ $\leq\frac{\mathrm{diam}(\Theta)^{2}+\beta^{2}L^{2}\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{2\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)}$
$=\frac{\beta L^{2}}{2}\frac{\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}L^{2}}+\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\,.$
By ˜J.1,
$\frac{\beta L^{2}}{2}\frac{\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}L^{2}}+\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}$ $\leq\frac{\beta L^{2}}{2}\frac{2\left(\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}L^{2}}+1+\log 2\right)}{\sqrt{T+2}}$
$=\frac{\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}L^{2}}{\beta}(T+2)^{-1/2}\,.$
Hence, ˜4.1 holds for $q=1/2$ as well. ∎
Remark J.6.
The proof idea for the case $q=1/2$ in ˜J.1 follows Beck (2017, Theorem 8.30).
J.2. PSGD (NSL)
The convergence rate for PSGD (NSL) is as follows.
Proposition J.7.
Let $\Theta$ be a bounded, closed, and convex set, and let $L>0$, $\theta^{1}\in\Theta$, and $T\in\mathbb{Z}_{\geq 1}$. Then, (A) for any $L$-Lipschitz convex function $\ell\colon\Theta\to\mathbb{R}$, the sequence $\{\theta^{t}\}_{t=1}^{T}$ produced by PSGD (NSL) satisfies
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\leq\frac{L}{2}\frac{\mathrm{diam}(\Theta)^{2}}{\sum_{t=1}^{T}\beta_{t}}+\frac{L}{2}\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\,.$
In particular, (B) there exists a function $Q_{L,\theta^{1}}\colon\mathbb{R}_{>0}\to\mathbb{R}$ satisfying ˜4.1.
proof.
(A) Substituting the NSL step size $\alpha_{t}=\beta_{t}/\|\nabla\ell(\theta^{t})\|$ into Equation˜J.3 yields
$\sum_{t=1}^{T}\frac{\beta_{t}}{\|\nabla\ell(\theta^{t})\|}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)\leq\frac{1}{2}\|\theta^{1}-\theta^{*}\|^{2}+\frac{1}{2}\sum_{t=1}^{T}\beta_{t}^{2}\,.$
On the other hand, since $\ell$ is $L$-Lipschitz, we have $\|\nabla\ell(\theta^{t})\|\leq L$, and thus
$\sum_{t=1}^{T}\frac{\beta_{t}}{\|\nabla\ell(\theta^{t})\|}\left(\ell(\theta^{t})-\min_{\Theta}\ell\right)$ $\geq\sum_{t=1}^{T}\frac{\beta_{t}}{\|\nabla\ell(\theta^{t})\|}\left(\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\right)$
$\geq\frac{1}{L}\left(\sum_{t=1}^{T}\beta_{t}\right)\left(\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\right).$
Combining the two displays proves (A).
(B) By the definition of NSL, we have $\sum_{t=1}^{\infty}\beta_{t}=\infty$. Hence, the right-hand side of (A) converges to 0 as $T\to\infty$, which implies the existence of $Q_{L,\theta^{1}}$ satisfying ˜4.1. ∎
Proposition J.8.
Consider the same setting as in ˜J.2. Let $0<q<1$ and $\beta>0$, and suppose that the step length is given by $\beta_{t}=\beta t^{-q}$. Define
$Q_{L,\theta^{1}}(\varepsilon)=\begin{cases}\left(\frac{L(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q))}{2\beta(1-q)\varepsilon}\right)^{1/q},&0<q<1/2,\\ \left(\frac{L(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q))}{2\beta(1-q)\varepsilon}\right)^{1/(1-q)},&1/2<q<1,\\ \max\left(1,\left(\frac{L(\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2})}{\beta\varepsilon}\right)^{2}-2\right),&q=1/2\,.\end{cases}$
Then $Q_{L,\theta^{1}}$ satisfies ˜4.1.
proof.
Suppose $0<q<1$ and $q\neq 1/2$. Proceeding as in the proof of ˜J.1, we obtain
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell\leq\begin{cases}\frac{L(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q))}{2\beta(1-q)}T^{-q},&0<q<1/2,\\ \frac{L(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q))}{2\beta(1-q)}T^{-(1-q)},&1/2<q<1.\end{cases}$
This implies the claim for $0<q<1$ with $q\neq 1/2$.
Now consider $q=1/2$. Again arguing as in the proof of ˜J.1, we have
$\min_{t=1,\ldots,T}\ell(\theta^{t})-\min_{\Theta}\ell$ $\leq\frac{L\mathrm{diam}(\Theta)^{2}+\beta^{2}L\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{2\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)}$
$=\frac{\beta L}{2}\frac{\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}}+\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\,.$ (J.6)
By ˜J.1,
$\frac{\beta L}{2}\frac{\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}}+\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}$ $\leq\frac{\beta L}{2}\frac{2\left(\frac{\mathrm{diam}(\Theta)^{2}}{\beta^{2}}+1+\log 2\right)}{\sqrt{T+2}}$
$=\frac{L}{\beta}\left(\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}\right)(T+2)^{-1/2}\,.$
Hence, the claim also holds for $q=1/2$. ∎
By Theorem˜4.3 and ˜J.2, we obtain the following corollary.
Corollary J.9.
Assume the setting of ˜J.2, and suppose that ˜3 and 3 hold. Then, if we implement Algorithm˜1 using PSGD under the present setting, the DDIOP Equation˜1.2 is solved exactly within at most
$\begin{cases}\left(\frac{L(\ell_{\mathrm{sub}})\left(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q)\right)}{2\beta(1-q)\gamma(\ell_{\mathrm{sub}})}\right)^{1/q},&0<q<1/2,\\ \left(\frac{L(\ell_{\mathrm{sub}})\left(\mathrm{diam}(\Theta)^{2}+\beta^{2}(1-2q)\right)}{2\beta(1-q)\gamma(\ell_{\mathrm{sub}})}\right)^{1/(1-q)},&1/2<q<1,\\ \max\left(1,\left(\frac{L(\ell_{\mathrm{sub}})\left(\mathrm{diam}(\Theta)^{2}+(1+\log 2)\beta^{2}\right)}{\beta\gamma(\ell_{\mathrm{sub}})}\right)^{2}-2\right),&q=1/2\end{cases}\,.$
iterations.
Appendix K Characterization of Piecewise Linear Mappings and Polyhedra
Definition K.1.
A bounded closed convex set $P$ is called a convex polytope if $P$ is the convex hull of a finite set. A bounded closed convex set $P$ is said to be a finite union of polyhedra if $P$ can be represented as the union of finitely many convex polytopes.
Definition K.2.
A mapping $f\colon\mathbb{R}^{m}\to\mathbb{R}$ is called an affine function if there exist $a_{1},\ldots,a_{m}\in\mathbb{R}$ and $b\in\mathbb{R}$ such that, for any $x=(x_{1},\ldots,x_{m})\in\mathbb{R}^{m}$,
$f(x)=a_{1}x_{1}+\cdots+a_{m}x_{m}+b$
holds. A mapping $f\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$ is called an affine map if every component $f_{i}$ of $f$ is an affine function.
Definition K.3.
Let $P\subset\mathbb{R}^{m}$ be a closed convex set. A function $f\colon P\to\mathbb{R}$ is said to be piecewise linear if there exists a finite family $\mathcal{Q}$ of convex polytopes such that
$P=\bigcup_{Q\in\mathcal{Q}}Q$
and, for every $Q\in\mathcal{Q}$, the restriction $f\Big\rvert_{Q}$ is an affine function. A mapping $f\colon P\to\mathbb{R}^{n}$ is said to be piecewise linear if every component $f_{i}$ of $f$ is piecewise linear.
Definition K.4.
Let $P$ be a closed convex set. A mapping $f\colon P\to\mathbb{R}^{n}$ is said to be polyhedral if its graph
$\Gamma(f)=\{(x,f(x))\mid x\in P\}$
is a finite union of polyhedra.
In this setting, piecewise linear functions can be characterized as follows.
Proposition K.5.
Assume that the domain $\operatorname{dom}f$ of a mapping $f$ is a convex polytope. Then the following statements are equivalent. (A) The mapping $f$ is polyhedral. (B) The mapping $f$ is piecewise linear. (C) The mapping $f$ can be represented as a composition of an affine map, $\min$, and $\max$.
It is also known that convex piecewise linear functions can be characterized as follows.
Proposition K.6 (Rockafellar 1970, Corollary 19.1.2, cf. Appendix˜K).
For a function $\ell\colon\Theta\to\mathbb{R}$, the following statements are equivalent. (A) The graph of $\ell$ is a convex polytope. (B) The function $\ell$ is convex and piecewise linear. (C) There exists a finite set $I$ such that, for any $i\in I$, there exist $a^{i}\in\mathbb{R}^{d}$ and $b_{i}\in\mathbb{R}$ satisfying
$\ell(\theta)=\max_{i\in I}\left(a^{i\top}\theta+b_{i}\right)$
In order to prove Appendix˜K, we first review definitions and propositions related to polyhedral mappings.
Definition K.7.
Let $\mathfrak{A}$ be the family of mappings generated by finite compositions of affine maps and
$\mathrm{min}_{2}(x_{1},x_{2},x_{3},\ldots,x_{m}):=(\min(x_{1},x_{2}),x_{3},\ldots,x_{m})$
Proposition K.8.
By Stallings and Swarup (1967, Proposition 1.6.9), the composition of two polyhedral mappings is polyhedral.
Proposition K.9.
By Stallings and Swarup (1967, Theorem 2.2.4), every polyhedral mapping is piecewise linear.
Proposition K.10.
By Ovchinnikov (2002, Theorem 2.1, §3), let $f$ be a piecewise linear mapping. Then each component $f_{i}$ of $f$ admits the representation
$f_{i}(x)=\min_{j\in J_{i}}\max_{k\in K_{ij}}g_{ijk}(x)\,$,
where $g_{ijk}$ are affine functions, and $J_{i}$ and $K_{ij}$ are finite sets.
Proof of Appendix˜K.
[…]
$f_{i}(x)=\min_{j\in J_{i}}\max_{k\in K_{ij}}g_{ijk}(x)$
, where $g_{ijk}$ are affine functions. On the other hand, for any $k\in\mathbb{Z}_{\geq 2}$,
$\mathrm{min}_{k+1}(x_{1},\ldots,x_{k+1},x_{k+2},\ldots,x_{m})$ $:=\mathrm{min}_{2}\Bigl(\mathrm{min}_{k}(x_{1},\ldots,x_{k}),x_{k+1},x_{k+2},\ldots,x_{m}\Bigr)\,$,
$\mathrm{max}_{k}(x_{1},\ldots,x_{k},x_{k+1},\ldots,x_{m})$ $:=-\mathrm{min}_{k}(-x_{1},\ldots,-x_{k},-x_{k+1},\ldots,-x_{m})$
belong to $\mathfrak{A}$. Therefore, $\min_{j\in J_{i}}$ and $\max_{k\in K_{ij}}$ can be represented as elements of $\mathfrak{A}$. Hence, $f\in\mathfrak{A}$.
[…]
It is immediate that affine maps are polyhedral. Next, we show that $\mathrm{min}_{2}$ is polyhedral. Let $P$ be a finite union of convex polytopes. Then there exists a finite family $\mathcal{Q}$ of convex polytopes such that
$P=\bigcup_{Q\in\mathcal{Q}}Q$
holds. The graph of $\mathrm{min}_{2}\Big\rvert_{P}$ is given by
$\Gamma\left(\mathrm{min}_{2}\Big\rvert_{P}\right)=\Gamma\left(\mathrm{min}_{2}\Big\rvert_{\bigcup_{Q\in\mathcal{Q}}Q}\right)=\bigcup_{Q\in\mathcal{Q}}\Gamma\left(\mathrm{min}_{2}\Big\rvert_{Q}\right)$
. For each $Q\in\mathcal{Q}$,
$\Gamma\left(\mathrm{min}_{2}\Big\rvert_{Q}\right)$
$=\{(x,\mathrm{min}_{2}(x))\mid x\in Q\}$
$=\{(x,\mathrm{min}_{2}(x))\mid x\in Q,\ x_{1}\geq x_{2}\}\cup\{(x,\mathrm{min}_{2}(x))\mid x\in Q,\ x_{2}\geq x_{1}\}$
$=\{(x,x_{1},x_{3},\ldots,x_{m})\mid x\in Q,\ x_{1}\geq x_{2}\}\cup\{(x,x_{2},x_{3},\ldots,x_{m})\mid x\in Q,\ x_{2}\geq x_{1}\}$
holds. Since affine maps send polyhedra to polyhedra, $\Gamma\left(\mathrm{min}_{2}\Big\rvert_{Q}\right)$ is a polyhedron. Hence, $\mathrm{min}_{2}$ is polyhedral. ∎
Proposition K.11.
Let $X^{\prime}\subset X\subset\mathbb{R}^{m}$ be polyhedra, and let $f\colon X\to\mathbb{R}^{d}$ be a polyhedral mapping. Then $f\Big\rvert_{X^{\prime}}$ is a polyhedral mapping.
proof.
Define the graph of the restriction $f\Big\rvert_{X^{\prime}}\colon X^{\prime}\to\mathbb{R}^{d}$ by
$\Gamma(f\Big\rvert_{X^{\prime}}):=\{(x,y)\in\mathbb{R}^{m+d}\mid x\in X^{\prime},\ y=f(x)\}$
Then,
$\Gamma(f\Big\rvert_{X^{\prime}})=\Gamma(f)\cap\bigl(X^{\prime}\times\mathbb{R}^{d}\bigr)$
holds. By assumption, $\Gamma(f)$ is a polyhedron; moreover, since $X^{\prime}$ is a polyhedron, $X^{\prime}\times\mathbb{R}^{d}$ is also a polyhedron. Therefore, since the intersection of polyhedra is a polyhedron, $\Gamma(f\Big\rvert_{X^{\prime}})$ is a polyhedron. ∎
Proposition K.12.
Let $X\subset\mathbb{R}^{m}$ be a polyhedron, and let $f\colon X\to\mathbb{R}^{d}$ be a polyhedral mapping. Then $f(X)\subset\mathbb{R}^{d}$ is a polyhedron.
proof.
Define the projection $\pi\colon\mathbb{R}^{m+d}\to\mathbb{R}^{d}$ by $\pi(x,y)=y$. By definition,
$f(X)=\{f(x)\mid x\in X\}=\{y\in\mathbb{R}^{d}\mid\exists x\in X,\ (x,y)\in\Gamma(f)\}=\pi(\Gamma(f))$
holds. Therefore, it suffices to show that if $\Gamma(f)$ is a polyhedron, then its linear image $\pi(\Gamma(f))$ is a polyhedron.
First, since $\Gamma(f)$ is a polyhedron, there exist finitely many convex polytopes $\{P_{j}\}_{j=1}^{J}$ such that
$\Gamma(f)=\bigcup_{j=1}^{J}P_{j}$
. Hence,
$f(X)=\pi(\Gamma(f))=\pi\Bigl(\bigcup_{j=1}^{J}P_{j}\Bigr)=\bigcup_{j=1}^{J}\pi(P_{j})$
and thus it suffices to show that each $\pi(P_{j})$ is a convex polytope.
Fix an arbitrary $j$ and set $P:=P_{j}$. Since $P$ is a convex polytope, there exist points $\{a^{i}\}_{i\in I}$ such that
$P=\mathrm{Conv}(\{a^{i}|i\in I\})$
. Since $\pi$ is linear,
$\pi(P)=\mathrm{Conv}(\{\pi(a^{i})|i\in I\})$
. Therefore, $\pi(P)$ is a convex polytope, and the proposition follows. ∎
Proposition K.13.
Let $X^{\prime}\subset X\subset\mathbb{R}^{m}$ be polyhedra, and let $f\colon X\to\mathbb{R}^{d}$ be a polyhedral mapping. Then $f(X^{\prime})$ is a polyhedron.
proof.
This follows from Appendixes˜K and K. ∎
Proposition K.14.
Let $X^{\prime}\subset X\subset\mathbb{R}^{m}$ be a finite union of polyhedra, and let $f\colon X\to\mathbb{R}^{d}$ be a polyhedral mapping. Then $f(X^{\prime})$ is a finite union of polyhedra.
proof.
Suppose that $X^{\prime}$ can be written, using a family of polyhedra $\{X^{i}\}_{i\in I}$, as
$X^{\prime}=\bigcup_{i\in I}X^{i}$
. Applying $f$ to both sides yields
$f(X^{\prime})=f\left(\bigcup_{i\in I}X^{i}\right)=\bigcup_{i\in I}f(X^{i})$
. By Appendix˜K, each $f(X^{i})$ is a polyhedron; hence $f(X^{\prime})$ is a finite union of polyhedra. ∎
Definition K.15.
Let $X$ be a finite union of polyhedra. A set $Y$ is said to be the vertex set of $X$ if $Y$ is the vertex set of the convex hull $\mathrm{Conv}X$.
Appendix L Technical Lemmas: Finite-time exact solvability of the DDIOP for MILPs
L.1. Sketch Proof of Theorem 4.3
In this section, we show that, for a convex piecewise-linear function whose set of minimizers has an interior point, one can attain the minimum value in finitely many iterations by employing a gradient-based optimization method that includes projected subgradient descent (Theorem˜L.3).
Assumption L.1.
Let $V\subset\mathbb{R}^{d}$ be an affine subspace. Let $\Theta\subset V$ be a convex body, i.e., a convex set with nonempty interior relative to $V$. Let $I$ be a finite set. For each $i\in I$, let $a^{i}\in\mathbb{R}^{d}$ and $b_{i}\in\mathbb{R}$. Assume that $\ell(\theta)=\max_{i\in I}\left(a^{i\top}\theta+b_{i}\right)$ and that $\operatorname*{arg\,min}_{\Theta}\ell$ has nonempty interior relative to $V$.
˜L.1 is equivalent to the condition that $\ell$ is convex and piecewise-linear and that its minimizer set has an interior point (cf. Appendix˜K). Under ˜L.1, the function $\ell$ is $\max_{i\in I}\|a^{i}\|$-Lipschitz continuous. Define $\widetilde{\ell}\colon\Theta\to\mathbb{R}$ and $\gamma(\ell)\in\mathbb{R}$ by
$\widetilde{\ell}(\theta):=\max_{i\in I\cap\{i\in I|a^{i}\not=0\}}\left(a^{i\top}\theta+b_{i}\right),\quad\gamma(\ell):=\min_{\Theta}\ell-\min_{\Theta}\widetilde{\ell}$ (L.1)
Remark L.2.
For a convex piecewise-linear function $\ell$, the definitions of $\widetilde{\ell}$ and $\gamma(\ell)$ are not necessarily well-defined in an intrinsic sense. Note that $\widetilde{\ell}$ is defined here only because $\ell$ is given in the explicit form $\ell(\theta)=\max_{i\in I}\left(a^{i\top}\theta+b_{i}\right)$.
Then, the following theorem holds.
Theorem L.3.
Suppose that ˜L.1, 4.1 and 4.1 hold. Let the Lipschitz constant be $L(\ell)=\max_{i\in I}\|a^{i}\|$. If $T\geq Q_{L(\ell),\theta^{1}}(\gamma(\ell))$, then $\min_{t=1,\ldots,T}\ell(\theta^{t})=\min_{\Theta}\ell$.
See ˜L.2 for the proof.
For the suboptimality loss, by the maximum principle, we can write
$\ell_{\mathrm{sub}}(\theta)=\frac{1}{N}\sum_{n=1}^{N}\max_{a^{n}\in Y^{(n)}}\theta^{\top}(a^{n}-a^{(n)})=\max_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}}\theta^{\top}\left(\frac{1}{N}\sum_{n=1}^{N}(a^{n}-a^{(n)})\right)$ (L.2)
The constant $\gamma(\ell_{\mathrm{sub}})$ is given as follows.
Proposition L.4.
Under ˜3 and 3, the constant $\gamma(\ell_{\mathrm{sub}})$ is given by Equation˜4.3.
See ˜L.4 for the proof.
Proof of Theorem˜4.3.
By ˜3 and 3, Equation˜L.2, and ˜5 and 3, the suboptimality loss satisfies ˜L.1. The theorem then follows from ˜L.1, ˜4.1 and 4.1, and Theorem˜L.3. ∎
L.2. A minimum-attainment theorem via gradient-based optimization for convex piecewise-linear functions
In this subsection, we prove Theorem˜L.3.
Proposition L.5.
Under ˜L.1, for any $\theta\in\Theta\setminus\operatorname*{arg\,min}_{\Theta}\ell$, the following hold: (A) $\ell(\theta)=\widetilde{\ell}(\theta)$; (B) $\partial\ell(\theta)=\partial\widetilde{\ell}(\theta)$.
proof.
(A) By definition, $\ell\geq\widetilde{\ell}$. Fix $\theta\in\Theta\setminus\operatorname*{arg\,min}_{\Theta}\ell$, and take $j\in\operatorname*{arg\,max}_{i\in I}(a^{i\top}\theta+b_{i})$. Since $\ell$ is convex, if $a^{j}=0$, then $\theta$ would attain the minimum value of $\ell$, which is a contradiction. Hence, $a^{j}\neq 0$. Therefore,
$\ell(\theta)$ $=\max_{i\in I}(a^{i\top}\theta+b_{i})=a^{j\top}\theta+b_{j}$
$=\max_{i\in I\cap\{i\mid a^{i}\neq 0\}}\left(a^{i\top}\theta+b_{i}\right)=\widetilde{\ell}(\theta).$
(B) Let $\mathrm{Conv}$ denote the convex hull. Then,
$\partial\ell(\theta)$ $=\mathrm{Conv}\left\{a^{i}\,\middle|\,a^{i\top}\theta+b_{i}=\max_{i\in I}(a^{i\top}\theta+b_{i})\right\},$
$\partial\widetilde{\ell}(\theta)$ $=\mathrm{Conv}\left\{a^{i}\,\middle|\,a^{i}\neq 0,\ a^{i\top}\theta+b_{i}=\max_{i\in I}(a^{i\top}\theta+b_{i})\right\}.$
By the same argument as in (A),
$\left\{a^{i}\,\middle|\,a^{i\top}\theta+b_{i}=\max_{i\in I}(a^{i\top}\theta+b_{i})\right\}=\left\{a^{i}\,\middle|\,a^{i}\neq 0,\ a^{i\top}\theta+b_{i}=\max_{i\in I}(a^{i\top}\theta+b_{i})\right\}.$
Hence, $\partial\ell(\theta)=\partial\widetilde{\ell}(\theta)$. ∎
Proposition L.6.
Under ˜L.1, one has $\gamma(\ell)>0$.
proof.
By definition, $\ell\geq\widetilde{\ell}$. Hence, $\gamma(\ell)\geq 0$.
Suppose, to the contrary, that $\gamma(\ell)=0$. Then, by definition, $\ell=\widetilde{\ell}$. For any $\theta\in\operatorname*{arg\,min}_{\Theta}\ell$, the subdifferential $\partial\ell(\theta)$ contains 0. Since $\operatorname*{arg\,min}_{\Theta}\ell$ has an interior point, at an interior point $\theta$ of $\operatorname*{arg\,min}_{\Theta}\ell$ we have $\partial\ell(\theta)=\{0\}$.
On the other hand, for any $\theta\in\Theta$, the subdifferential $\partial\widetilde{\ell}(\theta)$ necessarily contains
$\{a^{i}\mid i\in I,\ a^{i}\neq 0\}.$
Therefore, at an interior point $\theta$ of $\operatorname*{arg\,min}_{\Theta}\ell$, $\partial\ell(\theta)=\{0\}\neq\partial\widetilde{\ell}(\theta)$, which contradicts $\ell=\widetilde{\ell}$. Hence, $\gamma(\ell)>0$. ∎
Proposition L.7.
Under ˜L.1,
$\operatorname*{arg\,min}_{\Theta}\ell=\left\{\theta\in\Theta\,\middle|\,\widetilde{\ell}(\theta)\leq\min_{\Theta}\widetilde{\ell}+\gamma(\ell)\right\}.$
Proof of Theorem˜L.3.
Assume that, for all $t=1,\ldots,T$,
$\min_{t=1,\ldots,T}\ell(\theta^{t})>\min_{\Theta}\ell.$ (L.3)
Step 1: By ˜L.2(A), for every $t=1,\ldots,T$,
$\ell(\theta^{t})=\widetilde{\ell}(\theta^{t}).$ (L.4)
Moreover, by ˜L.2(B), we may identify
$\nabla\widetilde{\ell}(\theta^{t})=\nabla\ell(\theta^{t}).$ (L.5)
Step 2: Define a sequence $\{\theta^{\prime t}\}_{t}$ inductively by setting, for $t=1,\ldots,T-1$,
[…]
We show that $\theta^{t}=\theta^{\prime t}$ for all $t=1,\ldots,T$.
For $t=1$, we have $\theta^{1}=\theta^{\prime 1}$ by definition. Assume that $\theta^{t^{\prime}}=\theta^{\prime t^{\prime}}$ holds for all $t^{\prime}\leq t$. By the definitions of $\theta^{t+1}$, $\theta^{\prime t+1}$, and $\mathrm{update}_{t}$,
$\theta^{t+1}$ $=\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t}\mid\ell,\nabla\ell\right)=\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t},\{\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t},\{\nabla\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t}\right),$
$\theta^{\prime t+1}$ $=\mathrm{update}_{t}\left(\{\theta^{\prime t^{\prime}}\}_{t^{\prime}=1}^{t},\{\widetilde{\ell}(\theta^{\prime t^{\prime}})\}_{t^{\prime}=1}^{t},\{\nabla\widetilde{\ell}(\theta^{\prime t^{\prime}})\}_{t^{\prime}=1}^{t}\right).$
By Equations˜L.4 and L.5 and ˜4.1,
$\mathrm{update}_{t}\left(\{\theta^{t^{\prime}}\}_{t^{\prime}=1}^{t},\{\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t},\{\nabla\ell(\theta^{t^{\prime}})\}_{t^{\prime}=1}^{t}\right)$
$=\mathrm{update}_{t}\left(\{\theta^{\prime t^{\prime}}\}_{t^{\prime}=1}^{t},\{\widetilde{\ell}(\theta^{\prime t^{\prime}})\}_{t^{\prime}=1}^{t},\{\nabla\widetilde{\ell}(\theta^{\prime t^{\prime}})\}_{t^{\prime}=1}^{t}\right).$
Hence, $\theta^{t+1}=\theta^{\prime t+1}$. By induction, $\theta^{t}=\theta^{\prime t}$ for all $t=1,\ldots,T$.
Step 3: The Lipschitz constant of $\widetilde{\ell}$ is bounded by $\max_{i}\|a^{i}\|=L(\ell)$. If $T\geq Q_{L(\ell),\theta^{1}}(\gamma(\ell))$, then by ˜4.1,
$\min_{t=1,\ldots,T}\widetilde{\ell}(\theta^{t})-\min_{\Theta}\widetilde{\ell}\leq\gamma(\ell).$
Hence, there exists $t\in\{1,\ldots,T\}$ such that
$\widetilde{\ell}(\theta^{t})\leq\min_{\Theta}\widetilde{\ell}+\gamma(\ell).$
By ˜L.2, there exists $t\in\{1,\ldots,T\}$ such that
$\ell(\theta^{t})=\min_{\Theta}\ell,$
which contradicts Equation˜L.3. Therefore, Theorem˜L.3 holds. ∎
L.3. Proof of Proposition 5
Proof of ˜5.
(A) By Equation˜L.2, ˜3 and 3, and Appendix˜K, the suboptimality loss is piecewise linear.
(B) By definition,
$\theta^{\top}a^{*}(\theta,s^{(n)})-\theta^{\top}a^{(n)}\geq 0.$
Hence, $\ell_{\mathrm{sub}}(\theta)\geq 0$. Moreover, since $a^{(n)}=a^{*}(\theta^{*},s^{(n)})$ for every $n=1,\ldots,N$, we have $\ell_{\mathrm{sub}}(\theta^{*})=0$. Therefore, $\min_{\Theta}\ell_{\mathrm{sub}}=0$.
(C) Fix $n\in\{1,\ldots,N\}$. Define the open set $\Theta(a^{(n)},Y^{(n)})$ by
$\Theta(a^{(n)},Y^{(n)})=\left\{\theta\in\Theta\,\middle|\,\forall a^{n}\in Y^{(n)}\setminus\{a^{(n)}\},\ \theta^{\top}a^{(n)}>\theta^{\top}a^{n}\right\}.$
By ˜3 and $Y^{(n)}\subset f(X(s^{(n)}))$,
$\theta^{*}\in\Theta(a^{(n)},Y^{(n)}).$ (L.6)
On the other hand, for any $\theta\in\Theta(a^{(n)},Y^{(n)})$, we have $a^{*}(\theta,s^{(n)})=a^{(n)}$, and hence $\ell_{\mathrm{sub}}(\theta)=0$. By Equation˜L.6,
$\theta^{*}\in\bigcap_{n=1}^{N}\Theta(a^{(n)},Y^{(n)})\subset\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{sub}}.$
Since each $\Theta(a^{(n)},Y^{(n)})$ is open, the set $\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{sub}}$ contains an interior point. ∎
L.4. Proof of Proposition L.1
Proof of ˜L.1.
We compute the constant $\gamma(\ell_{\mathrm{sub}})$. Suppose that
$\sum_{n=1}^{N}\left(a^{n}-a^{(n)}\right)=0.$
Applying $\theta^{*\top}$ to both sides yields
$\sum_{n=1}^{N}\theta^{*\top}\left(a^{n}-a^{(n)}\right)=0.$
By the definition of $a^{(n)}$, we have $\theta^{*\top}\left(a^{n}-a^{(n)}\right)\leq 0$. Hence, for every $n=1,\ldots,N$,
$\theta^{*\top}\left(a^{n}-a^{(n)}\right)=0.$
By ˜3, it follows that
$a^{n}=a^{(n)}.$
Therefore,
$\sum_{n=1}^{N}\left(a^{n}-a^{(n)}\right)=0\Longleftrightarrow\forall n=1,\ldots,N,\ a^{n}=a^{(n)}.$
Consequently, the function $\widetilde{\ell_{\mathrm{sub}}}$ is given by
$\widetilde{\ell_{\mathrm{sub}}}(\theta)=\max_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{\top}\left(a^{n}-a^{(n)}\right).$ (L.7)
The claim follows from the definition of $\gamma(\ell_{\mathrm{sub}})$. ∎
Appendix M Universal finite-time exact solvability of the DDIOP for MILPs
As a technical condition, we introduce the following assumption.
Assumption M.1.
For almost every $\theta\in\Theta$, and for every $n=1,\ldots,N$, the feature $a^{*}(\theta,s^{(n)})$ is uniquely determined.
Example M.2.
By ˜3, the choice $\Theta=\Delta^{d-1}$ satisfies Appendix˜M.
Under Appendix˜M, define
$\Theta_{\mathrm{unique}}$
[…]
For any $n=1,\ldots,N$ and any $\theta\in\Theta_{\mathrm{unique}}$, condition Equation˜4.1 holds. Since each $Y^{(n)}$ is finite,
$\left\{(a^{*}(\theta,s^{(n)}))_{n}\,\middle|\,\theta\in\Theta_{\mathrm{unique}}\right\}\left(\subset\prod_{n=1}^{N}Y^{(n)}\right)$
is a finite set. By ˜L.1 and L.2, for any $\theta^{*}\in\Theta_{\mathrm{unique}}$, the following constants $\gamma_{\theta^{*}}$ and $\widetilde{\gamma}$ are strictly positive:
$\gamma_{\theta^{*}}$ $:=\max_{\theta\in\Theta}\min_{(a^{*}(\theta^{*},s^{(n)}))_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{\top}\left(a^{(n)}-a^{*}(\theta^{*},s^{(n)})\right),$
$\widetilde{\gamma}$ $:=\min_{\theta^{*}\in\Theta_{\mathrm{unique}}}\gamma_{\theta^{*}}.$ (M.1)
By definition, $\widetilde{\gamma}$ depends only on $\{X(s^{(n)})\}_{n}$, $f$, and $\Theta$.
Theorem M.3.
[…]
$\min_{t=1,\ldots,T}\ell_{\mathrm{sub}}(\theta^{t})=\min_{\Theta}\ell_{\mathrm{sub}}=0.$
proof.
[…]
$\min_{t=1,\ldots,T}\ell_{\mathrm{sub}}(\theta^{t})=0.$
Since Appendix˜M implies that $\Theta_{\mathrm{unique}}$ has full measure in $\Theta$, the conclusion holds for almost every $\theta^{*}\in\Theta$. ∎
Appendix N Technical Lemmas: finite-time attainment of the PLF minimum
We show that PSGD can drive a PLF to 0 within finitely many iterations.
Let $\theta^{**}\in\Theta$ satisfy
$\theta^{**}\in\operatorname*{arg\,min}_{\Theta}\widetilde{\ell_{\mathrm{sub}}}\left(\subset\operatorname*{arg\,min}_{\Theta}\ell_{\mathrm{sub}}\right).$
Lemma N.1.
Assume ˜3 and 3. Then, for any $\theta\in\Delta^{d-1}$, the following are equivalent: (A) $g(\theta)=0$; (B) $\ell_{\mathrm{plf}}(\theta)=0$.
Remark N.2.
Assume that the feature $a^{*}(\theta,s^{(n)})$ is defined as the lexicographically minimal element of
$\operatorname*{arg\,max}_{f(x)\in f(X(s))}\ \theta^{\top}f(x).$
Under this convention, Kitaoka and Eto (2023b) argued that once learning via Algorithm˜1 is completed, i.e., once a subgradient of the suboptimality loss becomes zero, one has $a^{*}(\theta,s^{(n)})=a^{(n)}$. The statement and proof of Appendix˜N are inspired by Kitaoka and Eto (2023b).
Proof of Appendix˜N.
(A) $\Rightarrow$ (B) Assume that a subgradient of $\ell_{\mathrm{sub}}$ is given by
$\sum_{n=1}^{N}\left(a^{*}(\theta,s^{(n)})-a^{*}(\theta^{*},s^{(n)})\right)=0.$
Applying $\theta^{*\top}$ to both sides yields
$\sum_{n=1}^{N}\theta^{*\top}a^{*}(\theta^{*},s^{(n)})=\sum_{n=1}^{N}\theta^{*\top}a^{*}(\theta,s^{(n)}).$
By the definition of $a^{*}(\theta^{*},s^{(n)})$, for every $n$,
$\theta^{*\top}a^{*}(\theta^{*},s^{(n)})\geq\theta^{*\top}a^{*}(\theta,s^{(n)}),$
and hence, for every $n$,
$\theta^{*\top}a^{*}(\theta^{*},s^{(n)})=\theta^{*\top}a^{*}(\theta,s^{(n)}).$
By Equation˜1.1,
$a^{*}(\theta,s^{(n)})\in\operatorname*{arg\,max}_{f(x)\in f(X(s^{(n)}))}\ \theta^{*\top}f(x).$
Since $\theta^{*}\in\Psi$, for every $n$ there exists $\xi^{(n)}\in Y^{(n)}$ such that
$\theta^{*}\in\Theta(\xi^{(n)},Y^{(n)}).$
This implies that, for every $n$, the set $\operatorname*{arg\,max}_{f(x)\in f(X(s^{(n)}))}\theta^{*\top}f(x)$ is a singleton. Therefore,
$a^{*}(\theta^{*},s^{(n)})=a^{*}(\theta,s^{(n)}).$
(B) $\Rightarrow$ (A) Under (B), for every $n$,
$a^{*}(\theta,s^{(n)})-a^{*}(\theta^{*},s^{(n)})=0.$
Summing over $n$ yields (A). ∎
Lemma N.3.
Assume ˜3 and 3. Then, for any $\theta\in\Theta$ satisfying $g(\theta)\neq 0$,
$\gamma(\ell_{\mathrm{sub}})=\max_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{**\top}\left(a^{(n)}-a^{n}\right)\geq\theta^{**\top}g(\theta)$
holds.
proof.
By ˜L.1 and the choice of $\theta^{**}$,
$-\gamma(\ell_{\mathrm{sub}})=-\min_{\Theta}\widetilde{\ell_{\mathrm{sub}}}$ $=\min_{\theta\in\Theta}\max_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{\top}\left(a^{n}-a^{(n)}\right)$
$=\max_{(a^{n})_{n}\in\prod_{n=1}^{N}Y^{(n)}\setminus\{(a^{(n)})_{n}\}}\frac{1}{N}\sum_{n=1}^{N}\theta^{**\top}\left(a^{(n)}-a^{n}\right).$
The claim follows from ˜3(C). ∎
Proposition N.4.
Assume ˜3 and 3. Let $\{\theta^{t}\}_{t}$ be the sequence obtained by implementing the update rule $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 with PSGD. Then, for any $t\in\mathbb{Z}_{\geq 1}$,
$\|\theta^{t+1}-\theta^{**}\|^{2}\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\alpha_{t}\gamma(\ell_{\mathrm{sub}})+\alpha_{t}^{2}\left\|g(\theta^{t})\right\|^{2}.$
proof.
By Appendix˜N and $\theta^{t\top}g(\theta^{t})\geq 0$, the distance between $\theta^{t+1}$ and $\theta^{**}$ can be bounded as follows:
$\|\theta^{t+1}-\theta^{**}\|^{2}$ $\leq\left\|\operatorname{Proj}_{\Theta}\left(\theta^{t}-\alpha_{t}g(\theta^{t})\right)-\theta^{**}\right\|^{2}$
$\leq\left\|\theta^{t}-\alpha_{t}g(\theta^{t})-\theta^{**}\right\|^{2}$
$=\left\|\theta^{t}-\theta^{**}\right\|^{2}+2\alpha_{t}(\theta^{**}-\theta^{t})^{\top}g(\theta^{t})+\alpha_{t}^{2}\left\|g(\theta^{t})\right\|^{2}$
$\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\alpha_{t}\gamma(\ell_{\mathrm{sub}})+\alpha_{t}^{2}\left\|g(\theta^{t})\right\|^{2}.$
N.1. The case of PSGD (NSS)
Theorem N.5.
Assume that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence generated by implementing the update rules $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 via PSGD (NSS). Then, for sufficiently large $T\in\mathbb{Z}_{\geq 1}$, one has $\ell_{\mathrm{plf}}(\theta^{T})=0$.
proof.
By Appendix˜N, the problem of minimizing the PLF can be reduced to the problem of minimizing the convex suboptimality loss. Assume that $g(\theta^{t})\neq 0$ for all $t=1,\ldots,T$. By Appendix˜N,
$\|\theta^{t+1}-\theta^{**}\|^{2}\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\beta_{t}\gamma(\ell_{\mathrm{sub}})+\beta_{t}^{2}L(\ell_{\mathrm{sub}})^{2}.$
Summing both sides over $t=1,\ldots,T$ yields
$\|\theta^{T+1}-\theta^{**}\|^{2}$ $\leq\left\|\theta^{1}-\theta^{**}\right\|^{2}-2\gamma(\ell_{\mathrm{sub}})\sum_{t=1}^{T}\beta_{t}+L(\ell_{\mathrm{sub}})^{2}\sum_{t=1}^{T}\beta_{t}^{2}$
$=\left\|\theta^{1}-\theta^{**}\right\|^{2}-\left(\sum_{t=1}^{T}\beta_{t}\right)\left(2\gamma(\ell_{\mathrm{sub}})-L(\ell_{\mathrm{sub}})^{2}\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\right).$ (N.1)
By the assumption on the sequence $\{\beta_{t}\}_{t}$, for sufficiently large $T\in\mathbb{Z}_{\geq 1}$,
$\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\leq\gamma(\ell_{\mathrm{sub}})$
holds, and thus Equation˜N.1 becomes strictly negative for sufficiently large $T$. This contradicts $\|\theta^{T+1}-\theta^{**}\|^{2}\geq 0$.
Therefore, there exists $t\in\{1,\ldots,T\}$ such that $g(\theta^{t})=0$. By Appendix˜N, there exists $t\in\{1,\ldots,T\}$ such that $\ell_{\mathrm{plf}}(\theta^{t})=0$. ∎
Theorem N.6.
Assume that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence generated by implementing the update rules $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 via PSGD (SRSS). Then, for any $T\in\mathbb{Z}_{\geq 1}$, either $\ell_{\mathrm{plf}}(\theta^{T})=0$, or
$\|\theta^{T+1}-\theta^{**}\|^{2}\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right).$
proof.
By Appendix˜N, the problem of minimizing the PLF can be reduced to the problem of minimizing the convex suboptimality loss. Assume that, for $t=1,\ldots,T$, one has $g(\theta^{t})\neq 0$. By Appendix˜N,
$\|\theta^{t+1}-\theta^{**}\|^{2}$ $\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\beta_{t}\gamma(\ell_{\mathrm{sub}})+\beta_{t}^{2}L(\ell_{\mathrm{sub}})^{2}$
$=\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\beta\gamma(\ell_{\mathrm{sub}})t^{-1/2}+\beta^{2}L(\ell_{\mathrm{sub}})^{2}t^{-1}.$
Summing over $t=\lceil T/2\rceil,\lceil T/2\rceil+1,\ldots,T$, we obtain
$\|\theta^{T+1}-\theta^{**}\|^{2}$
$\leq\left\|\theta^{\lceil T/2\rceil}-\theta^{**}\right\|^{2}-2\beta\gamma(\ell_{\mathrm{sub}})\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}+\beta^{2}L(\ell_{\mathrm{sub}})^{2}\sum_{t=\lceil T/2\rceil}^{T}t^{-1}$
$\leq\left\|\theta^{\lceil T/2\rceil}-\theta^{**}\right\|^{2}-\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\right)$
$\leq\mathrm{diam}(\Theta)^{2}-\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\right).$
By ˜J.1,
$\frac{\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}$ $\leq\frac{2(1+\log 2)}{\sqrt{T+2}}.$
Therefore,
$\mathrm{diam}(\Theta)^{2}-\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{\sum_{t=\lceil T/2\rceil}^{T}t^{-1}}{\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}}\right)$
$\leq\mathrm{diam}(\Theta)^{2}-\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right).$
Moreover,
$\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}$ $\geq\sum_{t=\lceil T/2\rceil+1}^{T}t^{-1/2}\geq\int_{\lceil T/2\rceil}^{T}z^{-1/2}dz=2T^{1/2}-2\lceil T/2\rceil^{1/2}$
$=2T^{1/2}-\sqrt{2}(T+1)^{1/2}=\frac{4T-2(T+1)}{2T^{1/2}+\sqrt{2}(T+1)^{1/2}}\geq\frac{2T-2}{(2+\sqrt{2})T^{1/2}}$
$=(2-\sqrt{2})\left(\sqrt{T}-1\right).$
Hence,
$\mathrm{diam}(\Theta)^{2}-\beta\left(\sum_{t=\lceil T/2\rceil}^{T}t^{-1/2}\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right)$
$\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right),$
which proves the claim. ∎
Proof of Theorem˜4.10.
(A) This follows from Theorem˜N.6.
(B) By the assumption on $T$,
$2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\geq\gamma(\ell_{\mathrm{sub}}).$
By Theorem˜N.6,
$\|\theta^{T+1}-\theta^{**}\|^{2}$
$\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\gamma(\ell_{\mathrm{sub}})-\beta L(\ell_{\mathrm{sub}})^{2}\frac{2(1+\log 2)}{\sqrt{T+2}}\right)$
$\leq\mathrm{diam}(\Theta)^{2}-\beta\gamma(\ell_{\mathrm{sub}})(2-\sqrt{2})\left(\sqrt{T}-1\right).$
By the assumption on $T$,
$\mathrm{diam}(\Theta)^{2}-\beta\gamma(\ell_{\mathrm{sub}})(2-\sqrt{2})\left(T^{1/2}-1\right)\leq 0.$
Combining the above inequalities yields $\|\theta^{T+1}-\theta^{**}\|^{2}<0$, which is a contradiction.
Therefore, there exists $t\in\{1,\ldots,T\}$ such that $g(\theta^{t})=0$. By Appendix˜N, there exists $t\in\{1,\ldots,T\}$ such that $\ell_{\mathrm{plf}}(\theta^{t})=0$. ∎
N.2. The case of PSGD (NSL)
Theorem N.7.
Assume that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence generated by implementing the update rules $\{\mathrm{update}_{t}\}_{t}$ in Algorithm˜1 via PSGD (NSL). Then, for sufficiently large $T\in\mathbb{Z}_{\geq 1}$, one has $\ell_{\mathrm{plf}}(\theta^{T})=0$.
proof.
By Appendix˜N, the problem of minimizing the PLF can be reduced to the problem of minimizing the convex suboptimality loss. Assume that $g(\theta^{t})\neq 0$ for all $t=1,\ldots,T$. By Appendix˜N,
$\|\theta^{t+1}-\theta^{**}\|^{2}\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\beta_{t}\frac{\gamma(\ell_{\mathrm{sub}})}{\|g(\theta^{t})\|}+\beta_{t}^{2}\leq\left\|\theta^{t}-\theta^{**}\right\|^{2}-2\beta_{t}\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}+\beta_{t}^{2}.$
Summing both sides over $t=1,\ldots,T$ yields
$\|\theta^{T+1}-\theta^{**}\|^{2}$ $\leq\left\|\theta^{1}-\theta^{**}\right\|^{2}-2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}\sum_{t=1}^{T}\beta_{t}+\sum_{t=1}^{T}\beta_{t}^{2}$
$=\left\|\theta^{1}-\theta^{**}\right\|^{2}-\left(\sum_{t=1}^{T}\beta_{t}\right)\left(2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\right).$ (N.2)
By the assumption on the sequence $\{\beta_{t}\}_{t}$, for sufficiently large $T\in\mathbb{Z}_{\geq 1}$,
$\frac{\sum_{t=1}^{T}\beta_{t}^{2}}{\sum_{t=1}^{T}\beta_{t}}\leq\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}.$
Hence, for sufficiently large $T$, the right-hand side of Equation˜N.2 becomes strictly negative, contradicting $\|\theta^{T+1}-\theta^{**}\|^{2}\geq 0$.
Therefore, there exists $t\in\{1,\ldots,T\}$ such that $g(\theta^{t})=0$. By Appendix˜N, there exists $t\in\{1,\ldots,T\}$ such that $\ell_{\mathrm{plf}}(\theta^{t})=0$. ∎
Theorem N.8.
Suppose that ˜3 and 3 hold. Let $\{\theta^{t}\}_{t}$ be the sequence generated by implementing the update rule $\{\mathrm{update}_{t}\}_{t}$ of Algorithm˜1 via PSGD (SRSL). Then, for any $T\in\mathbb{Z}_{\geq 1}$, either $\ell_{\mathrm{plf}}(\theta^{T})=0$, or
$\|\theta^{T+1}-\theta^{**}\|^{2}\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\beta\frac{2(1+\log 2)}{\sqrt{T+2}}\right).$
proof.
The claim can be proved in the same manner as Theorem˜N.6. ∎
Proof of Theorem˜4.11.
(A) This follows immediately from Theorem˜N.8.
(B) By the assumption on $T$, we have
$2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\beta\frac{2(1+\log 2)}{\sqrt{T+2}}\geq\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}.$
Hence, by Theorem˜N.8,
$\|\theta^{T+1}-\theta^{**}\|^{2}$ $\leq\mathrm{diam}(\Theta)^{2}-\beta(2-\sqrt{2})\left(\sqrt{T}-1\right)\left(2\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}-\beta\frac{2(1+\log 2)}{\sqrt{T+2}}\right)$
$\leq\mathrm{diam}(\Theta)^{2}-\beta\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}(2-\sqrt{2})\left(\sqrt{T}-1\right).$
By the assumption on $T$, it follows that
$\mathrm{diam}(\Theta)^{2}-\beta\frac{\gamma(\ell_{\mathrm{sub}})}{L(\ell_{\mathrm{sub}})}(2-\sqrt{2})\left(T^{1/2}-1\right)\leq 0.$
Combining these inequalities yields $\|\theta^{T+1}-\theta^{**}\|^{2}<0$, which is a contradiction.
Therefore, there exists some $t\in\{1,\ldots,T\}$ such that $g(\theta^{t})=0$. By Appendix˜N, there exists some $t\in\{1,\ldots,T\}$ such that $\ell_{\mathrm{plf}}(\theta^{t})=0$. ∎
N.3. Some Remarks
Remark N.9.
Let $\ell\colon\mathbb{R}^{d}\to\mathbb{R}$ be an $L$-Lipschitz convex function. nemirovskiĭ1983problem established matching (tight) lower and upper bounds on the worst-case complexity of (deterministic) gradient-based first-order methods for ensuring that the function-value suboptimality is at most $\varepsilon>0$ in the problem
$\min_{x\in\mathbb{R}^{d},\ \|x\|\leq B}\ell(x),$
namely, $\Theta(L^{2}B^{2}/\varepsilon^{2})$ iterations.
By ˜J.2 and J.1, PSGD (SRSS, SRSL) attains this rate with respect to $\varepsilon$.
N.4. Remark: SPO loss
We define the estimation loss, namely the SPO loss (Mohajerin Esfahani et al., 2018; Elmachtoub and Grigas, 2022), by
$\ell_{\mathrm{est}}(\theta;\theta^{*}):=\frac{1}{N}\sum_{n=1}^{N}\left(\theta^{*\top}a^{(n)}-\theta^{*\top}a^{*}(\theta,s^{(n)})\right).$
Proposition N.10.
Assume that ˜3 and 3 hold. Then, for any $\theta\in\Theta$ satisfying $g(\theta)\neq 0$,
$\gamma(\ell_{\mathrm{sub}})\leq\ell_{\mathrm{est}}(\theta;\theta^{**}).$
proof.
By Appendix˜N,
$-\gamma(\ell_{\mathrm{sub}})\geq\theta^{**\top}g(\theta)=-\ell_{\mathrm{est}}(\theta;\theta^{**}),$
which implies the claim. ∎
Appendix O Experiments
O.1. Implementation and device details
The primary libraries used in our experiments are OR-Tools v9.8 (Perron and Furnon, 2023), NumPy 1.26.3 (Harris et al., 2020), and Python 3.9.0 (Van Rossum and Drake, 2009). All computations were conducted on a machine equipped with 192 Intel CPU cores and 1.0 TB of RAM.
O.2. Linear programming
Theoretical setup
We set $d\in\{4,6,8\}$, $n=N=1$, $J^{(n)}=100$, and $r_{\max}=10$. For each $n=1,\ldots,N$, we draw $\log_{0.1}r_{i}^{(n)}$ i.i.d. from the uniform distribution on $0,\log_{0.1}(1/r_{\max})]$. For each $n=1,\ldots,N$ and $j=1,\ldots,J^{(n)}$, we draw $b^{(n,j)}$ i.i.d. from the uniform distribution on $\mathbb{R}_{\geq 0}^{d}$. We then rescale each $b^{(n,j)}$ by a positive scalar so that $\sum_{i=1}^{d}r_{i}^{(n)2}b_{i}^{(n,j)2}=1$. Under this setting, we consider the following LP:
$\begin{split}\text{maximize}_{x}\,&\theta^{\top}x\\ \text{subject to}\,&\sum_{i=1}^{d}r_{i}^{(n)2}b_{i}^{(n,j)}x_{i}\leq 1,\\ &\forall j=1,\ldots,J^{(n)},\,x\geq 0.\end{split}$ (O.1)
Let $\mathcal{S}$ denote the set of pairs $(r,b)$.
CHAN setup
We apply the discussion in Chan et al. (2023, (15)) to Equation˜O.1. We rewrite Equation˜O.1 as
$\begin{split}\text{maximize}_{x}\,&\theta^{\top}x\\ \text{subject to}\,&A^{(n)}x\leq 1_{J^{(n)}}:=(\underbrace{1,\ldots,1}_{J^{(n)}\text{ times}}),\quad x\geq 0.\end{split}$
Applying the discussion in Chan et al. (2023, (15)) to Equation˜O.1, we obtain
$\begin{split}\text{minimize}_{x,\theta\in\Theta_{k}}\,&\sum_{n=1}^{N}\|x^{(n)}-a^{(n)}\|_{2}^{2}\\ \text{subject to}\,&A^{(n)\top}\lambda^{(n)}\geq\theta,\,\lambda^{(n)}\geq 0,\,\theta^{\top}x^{(n)}=1_{J^{(n)}}^{\top}\lambda^{(n)},\\ &A^{(n)}x^{(n)}\leq 1_{J^{(n)}},\,x^{(n)}\geq 0,\,\forall n=1,\ldots,N\,.\end{split}$ (O.2)
Results
We report the experimental results here. Figure˜5 shows, for $d=4,6,8$, the worst-case behavior of $\ell_{\mathrm{sub}}(\theta^{t}(\mathcal{D}))+0.001$, respectively.
Figure 5. Worst-case behavior of $\ell_{\mathrm{sub}}(\theta^{t}(\mathcal{D}))+0.001$ for the LP setting.
O.3. Single-machine weighted completion-time scheduling
Setup
We consider the single-machine weighted completion-time scheduling problem $1\mid r_{j}\mid\sum_{j}\theta_{j}C_{j}$, in which $d$ jobs are to be processed on a single machine. We assume that the machine can process at most one job at a time and that processing is non-preemptive. Let $j=1,\ldots,d$ index the jobs, and let $J=\{1,\ldots,d\}$ denote the job set. For each job $j$, let $p_{j}$ be its processing time, $\theta_{j}$ its importance weight, and $r_{j}$ its release time (i.e., the earliest time at which job $j$ can start). The goal is to find an execution order (schedule) on the machine so as to minimize the weighted sum of completion times, where $C_{j}$ denotes the completion time of job $j$.
Table 6. Decision variables and constraints in the scheduling formulation.
$d$ 4 6 8
Decision variables 16 36 64
Constraints 28 66 120
Let the continuous variable $b_{j}$ denote the start time of job $j$, and let $x_{jk}$ be a binary variable that equals $1$ if job $j$ precedes job $k$, and equals 0 otherwise. Define $M(p,r):=\max_{j}r_{j}+\sum_{j}p_{j}$. Then, the problem can be formulated as follows:
$\text{minimize}_{b,x}\,$ $\sum_{j\in J}\theta_{j}(b_{j}+p_{j})$
subject to $b_{j}+p_{j}-M(p,r)(1-x_{jk})\leq b_{k},\quad\forall j\neq k,$
$x_{jk}+x_{kj}=1,\ \ x_{jk}\in\{0,1\},\quad\forall j\neq k,$
$b_{j}\geq r_{j},\ \ b_{j}\in\mathbb{R},\quad\forall j.$
The numbers of decision variables and constraints are summarized in Table˜6.
We set
$\Theta=\left\{\theta\in\mathbb{R}^{d}\,\middle|\,\theta_{j}\geq 0,\,\sum_{j\in J}\theta_{j}=1\right\}+10^{-3}(1,\ldots,1)\,.$
We sample $r_{j}$ i.i.d. from the uniform distribution on $[0,10]$ and $p_{j}$ i.i.d. from the uniform distribution on $[1,5]$. Let $\mathcal{S}$ denote the set of pairs $s=(p,r)$, and let $X(s)$ denote the feasible set of $(b,x)$ satisfying the constraints. We set $N=1$. We run Algorithm˜1 on $\Theta$ in place of $\Delta^{d-1}$.
Results
We report the experimental results here. Figure˜6 shows, for $d=4,6,8$, the worst-case behavior of $\ell_{\mathrm{sub}}(\theta^{t}(\mathcal{D}))+0.001$, respectively.
Figure 6. Worst-case behavior of $\ell_{\mathrm{sub}}(\theta^{t}(\mathcal{D}))+0.001$ for the single-machine scheduling problem.
Table 7. Runtime of each method for $d=4$ jobs.
PSGD PSGD UPA RPA
(SRSL) (Polyak)
Mean (s) 0.010 3.70 9.12 2.06
Max (s) 0.024 103.49 12.17 2.86
Median (s) 0.009 0.024 8.74 1.98
Table 8. Runtime of each method for $d=6$ jobs.
PSGD PSGD UPA RPA
(SRSL) (Polyak)
Mean (s) 0.04 18.30 4.89 4.61
Max (s) 1.14 132.86 6.06 11.15
Median (s) 0.02 0.27 4.84 4.19
Table 9. Runtime of each method for $d=8$ jobs.
PSGD PSGD UPA RPA
(SRSL) (Polyak)
Mean (s) 0.382 86.72 7.52 33.83
Max (s) 2.27 1354.91 8.61 93.85
Median (s) 0.211 12.25 7.51 29.24